### v1

# Zeroth-Order Adaptive Neuron Alignment Based Pruning without Re-Training

Abstract
Network pruning is a set of computational techniques that aim to reduce a given model’s computational cost by removing a subset of its parameters while having minimal impact on performance. Throughout the last decade, the most widely used pruning paradigm has focused on pruning and re-training, which nowadays is inconvenient due to the vast amount of pre-trained models, which are in any case too expensive to re-train. In this paper, we exploit functional information from dense pre-trained models, i.e., their activations, to obtain sparse models that maximize the activations’ alignment w.r.t. their corresponding dense models. Hence, we propose NeuronAl, a top-up algorithm that can be used on top of any given pruning algorithm for LLMs, that modifies the block-wise and row-wise sparsity ratios to maximize the neuron alignment among activations. Moreover, differently from existing methods, our approach adaptively selects the best parameters for the block-wise and row-wise sparsity ratios w.r.t. to the model and the desired sparsity (given as input), and requires no re-training. We test our method on 4 different LLM families and 3 different sparsity ratios, showing how it consistently outperforms the latest state-of-the-art techniques. The code is available at https://github.com/eliacunegatti/NeuroAL.
1 Introduction
In recent times, Large Language Models (LLMs) have shown incredible performance over almost any language task [44, 33, 4]. However, their performance tends to grow with their sizes (i.e., the number of trainable parameters), which in turn is proportional to the computational burden required to train and then use such models. One way to reduce the computational cost of LLMs is through network pruning, i.e., algorithms that remove parameters while minimizing performance degradation. This approach has been extensively studied on Convolutional Neural Networks (CNNs) [13, 25, 43, 11], but nowadays the focus has shifted towards pre-trained models [41, 42, 22]. This shift has required a change of paradigm in pruning techniques: in fact, while in CNNs the main paradigm is iterative pruning (with re-training) [13], with pre-trained models (such as LLMs) in most cases it is not possible to fully re-train such models, because (1) training data are often not accessible, and (2) full re-training would be anyway too expensive. This calls for “exploiting” as much as possible the information contained in a pre-trained model to obtain a performant sparse version of it, using weight’s information [21], activations [40, 39], or reconstruction error [15], without the need for re-training. More recently, a new category of pruning algorithms, which we may call top-up algorithms (i.e., methods that can be applied on top of a given pruning algorithm for LLMs), has emerged, aiming at further improving pruning performance. Such approaches can be divided into two categories: those that minimize the reconstruction error [18, 45, 49], and those that impose non-uniform sparsity distribution modifying the block-wise sparsity [46]. The latter category is extremely effective for improving performance in CNNs [14, 38], while its application to LLMs is still limited [46].
Contributions In this paper, we first analyze the reasons behind the effectiveness of non-uniform sparsity distribution in sparse LLM. To do so, we carefully analyze the state-of-the-art top-up method OWL [46], investigating the reason underlying its better performance (w.r.t. other methods from the state of the art) as well as its limitations in terms of sensitiveness to its hyperparameters. Leveraging this knowledge, we introduce a new top-up method, that we call NeuronAl. The algorithm consists of a two-step approach that re-distributes the block-wise sparsity, i.e., the sparsity among Transformer blocks, and the row-wise sparsity, i.e., the sparsity for each row of a given layer’s matrix, maximizing the neuron alignment between dense and sparse activations. Contrary to OWL, NeuronAl does not require the user to specify any hyperparameter, as it automatically selects the most-performing values from a suitable set, hence adapting to the underlying model and the desired sparsity. Moreover, the use of alignment overcomes the necessity to define outliers on the activations, as done in OWL, which requires defining a threshold a priori. Finally, it does not use any gradient information [18, 45], hence significantly saving computational resources.
We test our approach on 3 Language Modeling datasets and 7 Zero-Shot tasks over 4 different LLM families from 7B to 13B parameters, to show its ability to outperform the most recent state-of-the-art techniques, including OWL [46] and DsNoT [49], over 3 different high sparsity values (60%, 70%, and 80%). To assess the robustness of our approach, we also conduct an in-depth ablation study.
2 Related Work
In this section, we provide a comprehensive discussion about network pruning applied to LLMs. We first introduce structured and unstructured network pruning; then, we focus on the latter, introducing the latest approaches proposed for improving sparse model performance.
Structured Network Pruning. Given a layer’s weight matrix $W\in\mathbb{R}^{n\times m}$ to sparsify, structured pruning removes either entire rows ($n$) or columns ($m$) (see the next section) aiming at speeding up both training and inference time. The first approach that applies structured pruning to LLMs has been proposed in [28], and focuses on the dependency of Transformers, i.e., it removes components of the networks while maximizing their original functionality. In [24], a pruning mechanism has been devised to remove components with the worst balance between loss and runtime. Other structured pruning approaches have been proposed based on combinatorial optimization [30], perturbative forward-pass only [9], and reduction of the embedding dimension through PCA [1]. Finally, in [17] it has been found that the last Transformer blocks are redundant, hence they can be completely removed with minor performance drops. The reason behind this phenomenon lies in the similarity between the learnable representation of consecutive blocks, which turns out to increase when the block depth increases. While all these approaches can achieve valuable inference speed-ups, the performance of the resulting sparse models w.r.t. their dense counterparts can be matched only at low sparsity values, such as 20% in [28] or 30% in [1]. This somehow limits the applicability of these methods, since in the case of models with billions of parameters one may need more aggressive pruning strategies to meet stringent hardware requirements.
Unstructured Network Pruning. Differently from structure pruning, unstructured pruning works by removing weights in a scattered (i.e., non-structured) way. While in this scenario the inference speed-up is limited (although techniques for reordering weights are available [26, 34, 50]), the performance w.r.t. the dense model can be preserved also at high sparsity ratios (i.e., above 50%), with the performance at lower sparsity being almost always completely preserved. The first approach of this kind has been proposed in [15], where weight pruning and reconstruction are combined based on the Hessian matrix. Even a simple magnitude-based approach turned out to perform well [21], as well when integrated with information on the neuron activations [40, 12]. All these approaches work by computing a score for each weight and then removing, uniformly for each layer, the worst-performing ones for a given sparsity ratio.
Top-Up Algorithms To improve the performance of unstructured pruning algorithms, several top-up algorithms have been devised. These approaches can be categorized into two distinct groups: methods that minimize the reconstruction error keeping the sparsity uniform for each block, and methods that modify the block-wise sparsity of the model resulting in non-uniform sparsity distribution across blocks.
The first group firstly sparsifies the model using a pruning algorithm and then, either dynamically [49] or by backpropagation [18], updates the pruning mask. The second group (to which our method belongs) modifies the block-wise sparsity (obtained by a given pruning algorithm) based either on activations’ outliers [46] or on layer-wise sparsity using block-wise reconstruction error with gradient information [45].
The idea of simply redistributing the layer-wise sparsity is known to be extremely well-performing on Multi-Layer Perceptrons (MLPs) and Convolutional Neural Networks (CNNs). The first approach of this kind, based on the Erdős–Rényi (ER) model, has been proposed in [35] for MLPs and then adjusted for CNNs in [11], while an empirical study about the effect of layer-wise pruning using different sparsity ratios has been done in [27]. Regarding Transformers (both for vision and text), the state-of-the-art algorithms [15, 40] have been devised to set the block-wise sparsity across the Transformer blocks in a uniform way. Later on, OWL has been proposed to build upon scoring-based pruning algorithms, adjusting the block-wise sparsity in a non-uniform way w.r.t. the number of outliers computed for each block. This approach improves the performance of several pruning algorithms, e.g. [15, 40], especially at sparsity above 60%. On the same line, BESA [45] allocates layer-wise sparsity across each block’s layer using gradient information. Recently, modality-wise sparsity distribution has been investigated in the case of multimodal tasks in [12, 19].
3 Observational Study
As discussed earlier, OWL [46] uses a non-uniform sparsity distribution across blocks, which provides an advantage in terms of performance w.r.t. uniform distribution. However, the reason why this occurs remains unclear. In the first part of this section, we focus on uncovering the reason behind this phenomenon. Furthermore, OWL suffers from some disadvantages: it requires the setting of two hyperparameters, one for the sparsity difference between consecutive blocks to utilize in the non-uniform distribution ($\lambda$), and one for the outlier definition ($M$), In the second part of this section, we investigate how the choice of the best hyperparameters is tied to the model and sparsity selected. This limitation leads to either selecting non-optimal hyperparameters or requiring a grid search to find the optimal values.
The observational study has been performed on two models (Phi-2 and LLama-1 7B) and three different sparsity ratios (0.6, 0.7, 0.8) using as pruning algorithm both Wanda [40] and multiflow [12] [^1] We did not include BESA [45] since it requires gradient information and works layer-based rather than block-based. We consider only gradient-free state-of-the-art methods..
[…]
We start by analyzing if the OWL’s performance improvement, compared to uniform sparsity distribution, could be unequivocally associated with its outlier-based score. Since the number of outliers for each block turns out to decrease with the layer depth, we test three straightforward non-uniform sparsity schedules (namely linear, exponential, and logarithmic), which do not require any forward step and do not depend on the outliers. Given a fixed parameter $\lambda$, these schedules work by redistributing the sparsity across layers in a monotonically increasing way (i.e., the sparsity of layer $i$ is always larger than the sparsity of layer $i-1,\forall i>1$). Fig. 1 displays the improvement, w.r.t. uniform distribution, achieved by the three sparsity schedules on two pruning algorithms (Wanda and multiflow) with $\lambda=0.08$ (as in [46]). The results highlight how non-uniform sparsity schedules, with no outlier information, can match, and in some cases even improve, the performance of OWL. Overall, the linear schedule turns out to be the most reliable one since it does not show oscillations in performance across the different sparsity ratios (while this happens for the logarithmic and exponential schedule).
Figure 1: Average perplexity improvement (percentage) w.r.t. uniform distribution, computed over 3 Language Modeling datasets when applying non-uniform distribution with linear, exponential, and logarithmic schedules. The horizontal dashed lines refer to the improvement w.r.t. uniform distribution achieved by OWL when applied to each of the two base pruning algorithms.
[…]
We then analyze the effect of the two hyperparameters set by OWL, namely $\lambda$ and $M$. The first hyperparameter is used to set how much the sparsity can vary across blocks (i.e., $[s-\lambda,s+\lambda]$) while keeping the overall sparsity fixed as $s$. In the experimentation reported in [46], $\lambda$ is set differently for each model, with $0.08$ being the most used value. As a matter of fact, we found that this value is not effective with all models and sparsity ratios. We test OWL, applied again to Wanda and multiflow, with 12 different values of $\lambda$, and computed the perplexity on the 3 Language Modeling datasets. As we show in Figure 2, it turns out that OWL achieves the best results for each model-sparsity combination for different values of $\lambda$, which suggests the difficulty of setting this hyperparameter a priori. The second hyperparameter, $M$, defines the outliers’ threshold: namely, for each block, the number of outliers is computed as the number of activations that are $M$ times greater than the block’s activations’ mean. We replicated the same analysis carried out for $\lambda$ and also for $M$, testing 12 different values on the 3 Language Modeling datasets when applying OWL to Wanda and multiflow. As for $\lambda$, we found that the optimal setting of $M$ heavily depends on the model and sparsity (noting that in [46] the most used value is $5$), as depicted in Figure 3.
Figure 2: Perplexity for different values of $\lambda$ over WikiText2 using OWL’s non-uniform distribution across blocks.
Figure 3: Perplexity for different values of $M$ over WikiText2 using $\lambda=0.08$.
[…]
4 Methodology
In this section, we propose our method, NeuronAl (Neuron Alignment). Given a pruning algorithm for a pre-trained LLM, our method re-computes the block-wise sparsity ratios for each Transformer block and the row-wise sparsity ratios based on the alignment between the dense activations and the sparse activations. The main strength of this method lies in its ability to be fully adaptive to the model and desired sparsity (given as input), requiring no single hyperparameter to be chosen a priori, but rather a set of suitable values from which NeuronAl can pick the most performing one.
Preliminaries Given a dense model $\mathcal{D}$, a pruning algorithm $\mathcal{P}$, and a desired sparsity $s$, unstructured network pruning generally computes a saliency score $\Psi$ for each weight $w\in\mathcal{D}$ and then binarizes these scores w.r.t. the $\mathtt{top}_{k}$ elements, where $k=1-|\mathcal{D}|\times s$. This allows to obtain a binary mask $\mathcal{M}$ to apply over $\mathcal{D}$, from which the final sparse model can be computed as $\mathcal{S}=\mathcal{D}\odot\mathcal{M}$. Since LLMs are composed of stacked Transformer blocks (each one denoted as $\mathcal{B}_{i}$), i.e., sets of linear layers (each one denoted as $\ell_{i}^{j}$) that implement the self-attention mechanism followed by an MLP, the binarization step is usually done uniformly per each layer $\ell_{i}^{j}$ [15, 40] as:
$\mathcal{M}^{\ell_{i}^{j}}=\mathtt{top}_{k^{\ell_{i}^{j}}}(\Psi^{\ell_{i}^{j}},\mathcal{D}^{\ell_{i}^{j}}).$ (1)
Neuron Alignment Our proposed method is based on the idea of combining the concept of neuron alignment, which requires no a priori definition of outliers (hence no $M$ parameter, as in OWL), with that of adaptivity, to remove the dependence from $\lambda$. Differently from the well-established reconstruction error [15], computed as $\arg\min_{M_{\ell},W_{\ell}}\left\|W_{\ell}X_{\ell}-(M_{\ell}\odot W_{\ell})X_{\ell}\right\|_{2}^{2}$ for each layer or block (hence measuring the difference between the sparse and dense representation of the output prior the non-linearity functions), neuron alignment provides a single scalar value that evaluates the model in its completeness focusing on the difference between sparse and dense activations, hence after the non-linearity. The method takes as input both $\mathcal{D}$ and its sparse version $\mathcal{S}$ generated by $\mathcal{P}$ with sparsity ratio $s$, and uses a small calibration data $C_{\lambda}$ to make a forward pass on both models, to retrieve the dense and sparse activations, respectively $\mathcal{A}_{\mathcal{D}}$ and $\mathcal{A}_{\mathcal{S}}$. The main idea behind NeuronAl is to maximize the neuron alignment by firstly modifying the vector of sparsity ratios for all blocks ($\mathbf{s}^{\mathcal{B}}$) and then for all rows ($\mathbf{s}^{r}$), where each row corresponds to the layer’s weight matrix $W^{\ell_{i}^{j}}$ (for each layer $\ell_{i}^{j}$ in $\mathcal{B}_{i}$), where $W^{\ell_{i}^{j}}\in\mathbb{R}^{r\times m}$. The main strength of this approach is that it does not require any weight update nor gradient information, but just a block- and row-wise recalibration and mask update via Eq. (1), using the same scoring criteria of $\mathcal{P}$.
[…]
4.1 Block-Wise Sparsity Ratio
The first step concerns the block-wise redistribution over the whole model. Our method takes as input the dense and sparse models ($\mathcal{D}$ and $\mathcal{S}$), the desired sparsity ratio ($s$), the calibration data $C_{\lambda}$, and a set of different $\lambda$ parameters ($\lambda^{\text{set}}$). Then, it computes a set of $|\lambda^{\text{set}}|$ different vectors of block-wise sparsity values for the whole model $\mathbf{s}^{\mathcal{B}}_{set}=\{\mathbf{s}^{\mathcal{B}}_{\lambda_{1}},\mathbf{s}^{\mathcal{B}}_{\lambda_{2}},\dots,\mathbf{s}^{\mathcal{B}}_{\lambda_{|\lambda^{\text{set}}|}}\}$, where each element $\mathbf{s}^{\mathcal{B}}_{\lambda_{k}}$ indicates a vector of block-wise sparsity values obtained with a linear schedule in $[s-\lambda_{k},s+\lambda_{k}]$. For each $\mathbf{s}^{\mathcal{B}}_{\lambda_{k}}$, we then forward the calibration data $C_{\lambda}$ through the model, and calculate the corresponding neuron alignment:
$neur_{al}=\sum_{\mathcal{B}_{i}}\sum_{\ell_{i}^{j}}\frac{\Bigg{\|}\tilde{A}_{D}^{j}-\tilde{A}_{{\mathcal{S}}_{\left(\mathbf{s}^{\mathcal{B}}_{\lambda_{k}}\right)}}^{j}\Bigg{\|}_{2}}{|\tilde{A}_{D}^{j}|}$ (2)
where $\tilde{A}$ means that the activations are normalized to sum up to one. Then, we select ${\left(\mathbf{s}^{\mathcal{B}}_{set}\right)^{*}}$, i.e., the $\lambda$ parameters per block that minimize Eq. (2). Finally, we update the block-wise sparsity with the selected ${\left(\mathbf{s}^{\mathcal{B}}_{set}\right)^{*}}$ via Eq. (1), thus obtaining a sparsified model $\mathcal{S}_{\mathcal{B}}$.
4.2 Row-Wise Sparsity Ratio
The second step is complementary to the previous one, but in this case, the sparsity is modified w.r.t. the rows of each layer. It is established [40] that pruning using the row as a comparison group [^2] Here, comparison group refers to the subset of weights whose scores are compared to decide which weights to prune. achieves better performance w.r.t. using the whole layer since it inherently maximizes the network connectivity [20, 7]. Here, we rely on such discovery to strengthen our approach and change the row-wise sparsity based on the neuron alignment of each layer. In this case, for each layer $\ell_{i}^{j}$ (i.e., for each $W^{\ell_{i}^{j}}\in\mathbb{R}^{r\times m}$) we redistribute the sparsity across the $r$ rows. Also in this case the $\lambda$ parameters are critical for deciding how to control the sparsity difference between consecutive rows. We take our sparse model obtained with the block-wise redistribution ($\mathcal{S}_{\mathcal{B}}$) and, for each layer $\ell_{i}^{j}$, we compute different row-wise sparsity values obtaining $\mathbf{s}^{s}_{set}=\{\mathbf{s}^{r}_{\lambda_{1}},\mathbf{s}^{r}_{\lambda_{2}},\dots,\mathbf{s}^{r}_{\lambda_{|\lambda^{\text{set}}|}}\}$, where each $\mathbf{s}^{r}_{\lambda_{k}}$ indicates a vector of row-wise sparsity ranging in $[s-\lambda_{k},s+\lambda_{k}]$, where each element is inversely proportional to the alignment of the corresponding row. In this case, we select in $\mathbf{s}^{s}_{set}$ the row-wise vector ${\left(\mathbf{s}^{r}_{set}\right)^{*}}$ that minimizes Eq. (2), and then apply Eq. (1) to $\mathcal{S}_{\mathcal{B}}$, using each row as comparison group.
The full procedure composed of both the block-wise and row-wise recalibration is summarized in Algorithm 1.
Algorithm 1 Proposed top-up pruning procedure
Require: $\mathcal{D},\mathcal{P},s,C_{\lambda},\lambda^{\text{set}}$
$\mathcal{M}\leftarrow\mathcal{P}(\mathcal{D},s)$
$\mathcal{S}\leftarrow\mathcal{D}\odot\mathcal{M}$triangleright Prune $\mathcal{D}$ uniformly per layer
$A_{\mathcal{D}}\leftarrow\mathcal{D}(C_{\lambda})$triangleright Dense activations
$\left(\mathbf{s}^{\mathcal{B}}_{set}\right)^{*}\leftarrow$ NeuronAl($\mathcal{D},\mathcal{S},\lambda^{\text{set}},C_{\lambda},A_{\mathcal{D}}$) $\triangleright$ Block-wise step
$\mathcal{S}_{\mathcal{B}}\leftarrow\mathcal{D}\odot\mathtt{top}_{\left(\mathbf{s}^{\mathcal{B}}_{set}\right)^{*}}(\Psi,\mathcal{D})$
${\left(\mathbf{s}^{r}_{set}\right)^{*}}\leftarrow\textsc{{NeuronAl}}(\mathcal{D},\mathcal{S}_{\mathcal{B}},\lambda^{\text{set}},C_{\lambda},A_{\mathcal{D}})$triangleright Row-wise step
$\mathcal{S}_{\text{final}}\leftarrow\mathcal{D}\odot\mathtt{top}_{\left(\mathbf{s}^{r}_{set}\right)^{*}}(\Psi,\mathcal{D})$
function NeuronAl($\mathcal{D},\mathcal{S},s,\lambda^{\text{set}},C_{\lambda},A_{\mathcal{D}}$)
     $\mathbf{s}^{*}\leftarrow\emptyset$, $neur_{al}^{*}\leftarrow\infty$
     for $\lambda\in\lambda^{\text{set}}$ do
         $\mathbf{s}_{\lambda}$ = GetDist($s$,$\lambda$) $\triangleright$ Compute $\mathbf{s}^{\mathcal{B}}_{\lambda}$ or $\mathbf{s}^{r}_{\lambda}$
         $A_{\mathcal{S}}\leftarrow(\mathcal{D}\odot\mathtt{top}_{\mathbf{s}_{\lambda}}(\Psi,\mathcal{D}))(C_{\lambda})$triangleright Sparse activations
         $neur_{al}\leftarrow\textsc{GetAlign}(A_{\mathcal{D}},A_{\mathcal{S}},\mathbf{s}_{\lambda})$triangleright Via Eq. (2)
         if $neur_{al}<neur_{al}^{*}$ then
              $\mathbf{s}^{*}\leftarrow\mathbf{s}_{\lambda}$
     return $\mathbf{s}^{*}$
5 Experiments
We apply our proposed NeuronAl to different pruning algorithms tailored for LLMs. Specifically, we test how it compares in terms of performance over Language Modeling datasets and Zero-Shot tasks w.r.t. the most recent top-up algorithms for pruning. We also perform scalability and ablation studies to show the effectiveness of our NeuronAl.
5.1 Experimental Setup
Language Modeling Datasets To measure the models’ perplexity on Language Modeling datasets, we use the following three datasets: (1) WikiText2 [31], a collection of $2.6M$ tokens from Wikipedia; (2) Colossal Clean Common Crawl (C4) [37], which is a clean version of the Common Crawl dataset produced in April 2019, containing $154G$ tokens [10]; and (3) Penn Treebank (PTB), which is a dataset extracted from the Wall Street Journal containing $2.4M$ tokens.
Zero-Shot Tasks To assess more thoroughly how the different pruning algorithms affect the models’ capabilities, we employ the following 7 datasets. (1) Recognizing Textual Entailment (RTE) [8, 2, 16, 3], a dataset composed of $3000$ test samples in which a model has to detect the entailment between two sentences. (2) WinoGrande [23], a dataset consisting of $1767$ test cases in which the model has to fill a blank using two choices. (3) BoolQ [5], a question-answering dataset containing $3270$ test samples, where each question requires a yes/no answer. (4) HellaSwag [47], a dataset composed of $10k$ test cases that requires a model to select the appropriate completion for a given sentence from a set of possibilities. (5) ARC-e [6], the “easy” fold from the AI2 Reasoning Challenge, containing $2376$ questions with multiple-choice answers. (6) ARC-c [6], the “challenge” fold from the AI2 Reasoning Challenge, containing $1172$ questions that were not solvable for methods that perform well on the ARC-e dataset. (7) OBQA [32], a dataset containing $500$ test multiple-choice questions about core science facts, asked in novel contexts.
Table 1: Perplexity on the 3 Language Modeling datasets computed over 5 different LLMs for four different top-up algorithms (Uniform, DSnoT, OWL, and NeuronAl) on 3 pruning algorithms (Magnitude, multiflow, and Wanda) at 70% sparsity.
Algorithm Top-Up Phi-2.7B LLama-1 7B LLama-2 7B Mistral 7B OPT 6.7B WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB Magnitude Uniform 764.6 384.4 983.9 2.53e4 2.25e4 3.26e4 1.42e5 1.02e4 2.02e6 221.9 232.9 748.7 1.00e4 5.39e3 6.54e3 DSnoT 539.0 258.0 656.2 1.02e7 2.77e6 4.99e7 1.31e8 2.90e7 2.25e8 192.7 189.9 566.2 6.16e3 3.93e3 4.36e3 OWL 419.6 242.7 358.5 1.20e4 6.58e3 5.39e4 3.39e5 1.24e4 3.28e6 111.7 124.2 545.5 1.57e4 8.48e3 9.67e3 NeuronAl 278.6 177.0 316.0 208.7 203.8 3.70e3 155.8 264.8 2.61e3 46.1 42.7 552.7 2.10e4 1.07e4 1.09e4 multiflow Uniform 388.4 298.8 610.8 80.9 71.9 172.4 60.0 58.8 1260.4 936.8 656.3 2062.5 943.6 1250.4 843.1 DSnoT 325.5 261.9 328.8 67.6 65.0 114.7 66.6 75.8 688.9 57.4 63.3 264.8 241.8 153.3 263.9 OWL 197.9 141.3 293.9 25.1 25.8 78.9 29.2 31.0 547.1 329.0 764.3 1718.2 240.9 495.6 337.8 NeuronAl 110.7 91.0 192.0 20.6 21.2 45.9 22.0 23.8 265.5 197.2 343.9 954.0 221.7 86.5 219.7 Wanda Uniform 227.6 182.7 346.2 85.1 86.2 157.0 78.0 81.0 599.3 60.7 73.6 298.3 157.5 260.1 209.2 DSnoT 221.9 172.6 257.6 72.9 76.0 121.0 76.1 85.7 491.8 81.3 79.9 304.8 191.4 173.3 182.6 OWL 132.7 116.2 183.7 24.6 27.3 61.2 30.5 36.6 333.7 41.0 51.8 253.5 54.4 69.7 100.7 NeuronAl 89.9 78.9 133.9 21.5 23.2 43.8 23.9 27.2 205.3 28.9 33.9 197.8 183.9 87.3 194.7
Baselines As pruning algorithms, we test Magnitude, multiflow [12], and Wanda [40]. All are tested with three different top-up algorithms (besides ours): (1) Uniform distribution, (2) DsnoT [49] (dynamic training-free uniform distribution with mask update), and (3) OWL [46] (block-wise training-free non-uniform distribution based on outliers scores). All these baselines are tested considering as comparison group each row: in other words, for a given layer, the sparsity $s$ is uniform for each row of each matrix rather than uniform across matrices. This is done for two main reasons: 1) as mentioned earlier, it is established that row-wise pruning on LLMs leads to better performance w.r.t. layer-wise pruning [40], and 2) since our approach relies on a row-wise step, for fairness we also use row as comparison group (rather than layer) on all the other methods, to appreciate the benefit of our approach. We also test our approach, as well as the baselines, on SparseGPT [15], using in this case only the block-wise approach, since SparseGPT relies on a weight reconstruction mechanism that firstly prunes columns and then updates the rows of the pruned cells, which makes it unfeasible to apply our row-wise approach. The results on SparseGPT can be found in Tables 14-16 in the Appendix. For all the pruning algorithms that use calibration data (i.e., multiflow, Wanda, and SparseGPT), we use 128 samples from the C4 dataset, as in [15, 40].
Models and Sparsity Since one of the main distinctive features of NeuronAl is its ability to adapt to sparsity and models, we test 4 different LLM families. Specifically, we evaluate LLama-7B (both V1 and V2) [41, 42], Phi-2, Mistral-7B [22], and OPT-6.7B [48]. To scale up the models’ size, we also test LLama-13 (both V1 and V2). In the paper, we mainly present results at 70% sparsity, for fair comparisons with [46]. However, to assess the generalization to different sparsity ratios we also include 60% and 80% sparsity in our experiments, see the Appendix for details.
NeuronAl Setup Our proposed method works by taking as input an LLM model, a desired sparsity ratio, a scoring-based pruning algorithm, and two sets of $\lambda$ parameters (one for the block-wise and one for the row-wise steps). In our experiments, we gave as input $\lambda^{\text{set}}\in[0.01,0.02,0.03,0.05,0.06,0.07,0.08,0.09,0.1,0.12,0.15$, 0.20,0.25]0.20,0.25]0.20 , 0.25 ] for the block-wise step, while for the row-wise step we also added $0.0$ (in case of no performance improvement). Each value in $\lambda^{\text{set}}$ has been evaluated (as described in Algorithm 1) over a calibration data $C_{\lambda}$. Since all the base pruning algorithms require a single forward over $C$ (where $|C|=128$) sequences with 2048 tokens each, while OWL requires a second forward always over $C$, to make the computational runtime similar we set $C_{\lambda}=8$ (the closest power of 2 w.r.t. $|C|/|\lambda^{\text{set}}|$) [^3] In Fig. 11 in the Appendix we report the results with different sizes of $C_{\lambda}$, showing how NeuronAl is robust to $|C_{\lambda}|$.. Doing so, the computational time required by NeuronAl is only two forward steps more than the base pruning algorithms, and one forward more than OWL [^4] For both $C$ and $C_{\lambda}$, we use the same seed (00) for the calibration set, i.e., $C_{\lambda}$ contains the first 8 elements of $C$.. All the experiments have been run on NVIDIA A100 GPUs, both with 40 and 80 GB.
5.2 Experimental Results
Concerning the Language Modeling datasets, the numerical results in terms of perplexity computed over the 3 Language Modeling datasets at 70% sparsity are shown in Table 1. It can be seen how NeuronAl is able in almost all cases to outperform all the other baselines by a large margin. In no case NeuronAl performs worse w.r.t. the uniform distribution. The only model on which NeuronAl is not the best top-up algorithm for all pruning algorithms is OPT. In all other cases, NeuronAl outperforms OWL for all models and pruning algorithms. The results at 60% and 80% sparsity shown in Tables 6-7 in the Appendix confirm this trend.
As for the Zero-Shot tasks, the numerical results are shown in Table 2555Magnitude is removed from this table due to space constraints. Its results are available separately in Tables 10-12 in the Appendix.. Again, NeuronAl turns out to outperform in the majority of cases all the baselines. In 8 cases out of 10 (w.r.t. the mean accuracy across all tasks), NeuronAl is the one that reaches the best performance. The results for 60% and 80% sparsity are available in Tables 8-9 in the Appendix, where we observe a similar trend.
Table 2: Zero-Shot accuracy for 7 tasks, computed over 5 different LLMs for 3 different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on two pruning algorithms (multiflow and Wanda) at 70% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B multiflow 53.79 50.43 53.49 28.43 46.59 20.48 15.6 38.40 w. DSnoT 52.71 54.06 54.92 28.38 44.15 21.5 15.6 38.76 w. OWL 52.35 55.33 61.99 30.31 54.59 24.23 18.4 42.46 w. NeuronAl 53.07 58.01 62.17 33.33 51.52 25.43 17.6 43.02 Wanda 52.35 53.2 62.14 28.31 44.87 20.99 17.4 39.89 w. DSnoT 52.35 51.54 60.98 28.33 41.62 21.08 15.4 38.76 w. OWL 52.71 53.59 62.05 30.09 48.61 22.53 18.8 41.20 w. NeuronAl 52.71 58.01 62.17 32.37 49.62 22.95 17.2 42.15 Llama1 7B multiflow 55.96 52.57 61.96 29.77 34.64 19.45 15.2 38.51 w. DSnoT 54.15 50.43 59.33 29.33 36.45 19.28 13.6 37.51 w. OWL 52.35 58.64 62.63 36.74 47.43 26.62 18.2 43.23 w. NeuronAl 58.12 61.33 63.55 38.23 50.76 26.71 22.6 45.90 Wanda 55.23 52.8 57.46 28.84 32.2 18.0 13.8 36.90 w. DSnoT 54.15 51.22 54.56 28.97 33.08 18.26 13.6 36.26 w. OWL 58.48 58.56 62.60 34.74 47.35 24.06 17.4 43.31 w. NeuronAl 54.51 58.72 63.30 37.06 49.92 26.37 20.6 44.35 Llama2 7B multiflow 52.71 50.99 62.05 28.52 33.04 17.92 13.6 36.98 w. DSnoT 52.71 50.99 59.72 27.92 32.58 16.81 13.0 36.25 w. OWL 52.71 56.12 62.05 32.40 42.42 19.88 18.6 40.60 w. NeuronAl 54.15 58.09 62.42 35.29 48.36 22.53 21.2 43.15 Wanda 52.71 48.46 49.94 28.09 30.39 19.2 11.8 34.37 w. DSnoT 52.71 50.36 47.77 27.67 30.6 17.32 12.2 34.09 w. OWL 52.71 55.96 62.11 31.86 43.73 20.65 17.0 40.57 w. NeuronAl 52.71 57.62 62.72 35.47 50.25 22.01 20.6 43.05 Mistral-7B multiflow 49.82 50.75 41.19 26.45 26.64 21.84 12.6 32.76 w. DSnoT 52.71 52.57 62.42 29.51 36.66 18.94 12.0 37.83 w. OWL 53.79 49.17 38.90 26.77 27.78 19.20 12.8 32.63 w. NeuronAl 52.35 50.83 37.98 27.24 27.74 18.34 13.8 32.61 Wanda 52.71 51.62 59.79 28.86 34.18 18.17 12.6 36.85 w. DSnoT 52.71 50.28 58.62 28.51 33.54 18.86 13.2 36.53 w. OWL 52.71 53.91 62.20 30.95 39.39 18.60 13.6 38.77 w. NeuronAl 52.71 60.22 62.20 34.69 44.32 20.73 15.6 41.50 OPT-6.7B multiflow 53.79 49.72 43.0 26.48 30.51 20.05 13.8 33.91 w. DSnoT 53.79 49.01 61.1 27.01 32.87 18.34 12.2 36.33 w. OWL 48.74 48.62 61.56 27.18 35.69 16.47 11.6 35.69 w. NeuronAl 49.10 50.43 62.17 31.37 40.28 22.01 17.0 38.91 Wanda 52.71 49.72 60.03 26.91 35.86 17.75 11.2 36.31 w. DSnoT 52.71 49.57 60.61 26.91 35.06 17.58 12.0 36.35 w. OWL 53.79 51.22 61.87 29.53 42.3 18.09 14.6 38.77 w. NeuronAl 50.90 51.30 62.17 30.68 39.77 21.76 14.6 38.74
Table 3: Perplexity of LLama-13B on the 3 Language Modeling datasets at 70% (top) and 80% sparsity (bottom).
Algorithm Top-Up Llama-1 13B LLama-2 13B WikiText2 C4 PTB WikiText2 C4 PTB Magnitude Uniform 5610.6 3682.9 20040.0 191.0 159.1 2784.1 DSnoT 7420.1 2264.0 48325.4 121.7 103.0 2331.2 OWL 279.4 438.9 11502.2 33.3 35.4 1267.2 NeuronAl 55.4 83.3 841.1 45.2 44.4 321.1 multiflow Uniform 49.4 45.3 277.8 144.3 112.4 623.2 DSnoT 46.2 48.9 240.4 45.8 54.2 611.5 OWL 16.6 17.7 132.2 54.0 56.2 426.6 NeuronAl 13.7 15.4 98.7 17.3 20.0 275.2 Wanda Uniform 54.4 55.3 309.2 45.7 56.2 571.0 DSnoT 47.8 54.2 248.6 46.6 57.7 555.5 OWL 16.3 18.9 147.6 18.0 21.8 315.1 NeuronAl 14.3 16.6 85.4 16.5 19.3 239.6 Magnitude Uniform 26696.2 26319.4 27180.9 12020.1 11161.9 10210.6 DSnoT 5591.8 4131.3 5732.3 5822.9 6510.2 10668.0 OWL 21985.6 19564.4 24511.7 5208.2 5058.5 10202.2 NeuronAl 133.4 183.8 1727.7 574.3 386.2 4169.1 multiflow Uniform 3711.2 1696.1 3588.4 4481.2 2409.8 5209.6 DSnoT 5368.4 2863.6 6286.3 1943.3 1669.5 5279.7 OWL 813.8 375.7 2138.8 1802.3 1012.2 4394.9 NeuronAl 61.8 65.7 414.9 323.1 259.6 1301.6 Wanda Uniform 3479.3 1959.8 3566.1 1124.6 870.5 5549.8 DSnoT 43684.1 24432.2 32162.1 4441.2 3957.9 4092.8 OWL 761.6 368.1 1929.8 248.0 204.2 2027.1 NeuronAl 64.0 68.5 413.8 116.5 103.2 814.8
Scalability Study To assess if the NeuronAl performance scales to bigger models, we apply it to LLama-13B (both V1 and V2) on the Language Modeling datasets. The results available in Table 3 (top) show how our approach is even more effective with larger models. For the 70% sparsity ratio, NeuronAl turns out to reach the best performance over both models, for all pruning algorithms. Impressively, the gap in performance w.r.t. the other baselines (in particular uniform distribution) is even more marked than in the smaller models tested previously. The improvement of NeuronAl over the baselines is even more evident at 80% sparsity, see Table 3 (bottom), where in several cases it outperforms OWL by 1-2 orders of magnitude. This is confirmed also at 60% sparsity, see Table 13 in the Appendix.
NeuronAl vs block-only vs row-only Since NeuronAl is based on two consecutive steps (for brevity, we call them here block and row), we test the performance of the block and row steps applied alone. This means that in the block-only setting the sparsity of each block one is given by our NeuronAl, while the row-wise sparsity for each layer is fixed to $s$. Instead, for the row-only setting, the sparsity for each block is fixed at $s$, while the row-wise sparsity for each layer is given by our neuron alignment mechanism. Table 4 shows the results of such ablation. The results confirm that NeuronAl (with both steps) achieves better performance in most of the cases, although this depends on the sparsity ratio (at lower sparsity, the block-only setting is the best one in 3 out of 5 cases). It is worth noticing that in such cases the gap w.r.t. NeuronAl is marginal and that NeuronAl always outperforms the uniform distribution. However, on high sparsity NeuronAl always performs best, with a large margin w.r.t. the second best (which is always the block-only setting). It is also clear how most of the performance improvement of NeuronAl comes from the block step, while the row step yields a moderate improvement, which however increases when increasing the sparsity ratio.
Table 4: Ablation study: perplexity achieved when using the block and row steps alone over the C4 dataset.
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Uniform 29.3 13.7 14.0 15.9 17.9 NeuronAl block-only 27.3 11.8 11.9 13.8 17.8 NeuronAl row-only 29.2 14.0 14.5 15.9 19.7 NeuronAl 27.2 12.0 12.0 13.8 19.1 70% Uniform 182.7 86.2 81.0 73.6 260.1 NeuronAl block-only 82.0 22.4 25.6 35.4 91.4 NeuronAl row-only 175.6 85.4 88.9 73.6 152.0 NeuronAl 78.9 23.2 27.2 34.0 87.3 80% Uniform 12422.9 3969.8 3117.3 277.7 2345.8 NeuronAl block-only 3177.3 296.2 313.8 181.8 530.6 NeuronAl row-only 16501.8 4597.5 824.9 277.7 2345.8 NeuronAl 2404.5 242.0 226.2 168.6 471.4
NeuronAl $\lambda$ Selection In this section, we report an analysis of the ability of NeuronAl to pick the best $\lambda$ parameters (i.e., the parameters for which the performance is the best one, hence the lowest value if computed over perplexity). To do this, we evaluate NeuronAl for all the $\lambda$ parameters (in the block-only scenario to simplify the visualization of results) over the 3 Language Modeling datasets. Fig. 4 reports the perplexity at 70% sparsity across different values of $\lambda$ (black dots connected by solid lines), while the dots highlighted in orange indicate the perplexity achieved with the $\lambda$ value selected by NeuronAl. These results highlight how NeuronAl, in the majority of the cases, can pick the optimal value of $\lambda$ both with data knowledge, as in the C4 dataset (from which the calibration data is sampled), as well as on unseen datasets such as WikiText2 and PTB. Fig.s 9-10 in the Appendix show the results at 60%-80% sparsity.
Figure 4: Perplexity over different values of $\lambda$ at 70% sparsity. The orange dot indicates the value selected by NeuronAl. The best $\lambda$ is retrieved by maximizing the neuron alignment computed over $C_{\lambda}$ on the C4 dataset, and then it is transferred to the other datasets. The red line corresponds to the perplexity using a uniform distribution, hence $\lambda=0$.
Calibration Data Since NeuronAl works by finding the best $\lambda$ parameters w.r.t. to the neuron alignment using a calibration sample of 8 samples, we test the results under different seeds for the calibration data. The results in Table 5 show the mean and standard deviation of perplexity at 70% sparsity over the 3 Language Modeling datasets (see Tables 20-20 in the Appendix for the same ablation at 60% and 80% sparsity). It is visible how the standard deviation is minimal w.r.t. the mean, apart from the OPT model where the standard deviation turns out to be higher. This dependence on the calibration data could explain the reason why, as discussed before, our proposed approach works better than the baselines on all other models but struggles with the OPT family. The ablation on the number of samples (see Fig. 11 in the Appendix) supports this hypothesis. In fact, OPT models are the only ones to show performance oscillation depending on the calibration data seed and especially the number of samples, which does not happen with all the other LLM families.
Table 5: Ablation study: perplexity achieved when using NeuronAl with different calibration data seeds (0, 16, 46) at 70% sparsity over the 3 Language Modeling datasets.
Dataset Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B WikiText2 085.5 $\pm$ 3.3 22.7 $\pm$ 2.1 023.6 $\pm$ 0.2 029.0 $\pm$ 0.1 137.4 $\pm$ 71.2 C4 076.3 $\pm$ 2.2 24.2 $\pm$ 1.5 027.0 $\pm$ 0.2 034.2 $\pm$ 0.3 075.9 $\pm$ 16.2 PTB 135.8 $\pm$ 4.8 46.5 $\pm$ 5.0 200.4 $\pm$ 3.6 205.0 $\pm$ 6.7 154.5 $\pm$ 63.6
6 Conclusion and Limitations
In this paper, we proposed NeuronAl, a new approach to prune LLMs based on the alignment between sparse and dense activations. Our approach has the main strength of requiring no outliers or gradient information. On top of that, our method is also adaptive, since it is designed to automatically select the best hyperparameters for the model, pruning algorithm, and desired sparsity. Throughout extensive experiments, we showed how our approach outperforms all the latest state-of-the-art methods both on Language Modeling datasets and Zero-Shot tasks, on 4 different LLM families and 3 sparsity ratios. We also included an extensive ablation study to show the robustness of our approach to the calibration data as well as the effect of its algorithmic steps.
In our view, the present version of NeuronAl has three main limitations. (1) With NeuronAl, it is not possible to make use of optimized structured-sparsity inference implementations (e.g, the NVIDIA N:M sparsity [36]): for a given sparsity, NeuronAl produces customized sparsity constraints for each layer in a block, and then for each row of each layer. Therefore, these implementations cannot be employed as they often require continuity in the sparsity of matrices. (2) NeuronAl requires two forward steps more than base pruning algorithms, and one more than the closest top-up competitor, OWL. (3) Finally, it requires the definition of $\lambda^{\text{set}}$. However, it is worth noticing that $|\lambda^{\text{set}}|$ as well as $|\mathcal{C}_{\lambda}|$ affect the computational cost of the forward step.
Appendix A Appendix
A.1 Observational Study
Here we include the results of the observational study on the OWL’s hyperparameters $\lambda$ and $M$ computed on the C4 dataset (Fig.s 5 and 7), and PTB (Fig.s 6 and 8). The results confirm the findings presented in the main text: there is no single hyperparameter selection (both $M$ and $\lambda$) that reaches the best performance across all different model-pruning-sparsity settings.
Figure 5: Perplexity for different $\lambda$ values over C4 using OWL’s non-uniform distribution across blocks.
Figure 6: Perplexity for different $\lambda$ values over PTB using OWL’s non-uniform distribution across blocks.
Figure 7: Perplexity for different $M$ values over C4 using $\lambda=0.08$.
Figure 8: Perplexity for different $M$ values over PTB using $\lambda=0.08$.
A.2 Additional Experiments
Here we include the results of the experiments that, due to space limits, we could not include in the main text. Specifically, we report: the results of NeuronAl over Language Modeling and Zero-Shot tasks at 60% and 80% sparsity; the results for the Zero-Shot tasks with Magnitude pruning; the results over LLama-13B at 70% sparsity; the results of NeuronAl (block-only) applied to SparseGPT [15].
A.2.1 Language Modeling at 60% and 80% sparsity
In Table 6-7, we report the results of NeuronAl over the 3 Language Modeling datasets (WikiText2, C4, and PTB) with the five different models considered in the main text, for 60% and 80% sparsity. In the first case, our approach turns out to be the best one in 23 out of 45 cases, while for 80% sparsity in 27 out of 45. It is interesting to notice how at medium sparsity (60%) all the top-up algorithms, including ours, provide similar results, while the improvement provided by NeuronAl at 80% (w.r.t. the top-up competitors) in some cases reaches a factor 5x (e.g., with Llama-1 7B for multiflow and Wanda).
Table 6: Perplexity on the 3 Language Modeling datasets computed over 5 different LLMs for four different top-up algorithms (Uniform, DSnoT, OWL, and NeuronAl) on 3 pruning algorithms (Magnitude, multiflow, and Wanda) at 60% sparsity.
Algorithm Top-Up Phi-2.7B LLama-1 7B LLama-2 7B Mistral 7B OPT 6.7B WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB Magnitude Uniform 51.3 45.9 66.9 152.4 159.8 3016.1 6894.8 42666.8 1706134.9 19.6 24.4 189.3 9489.3 6203.7 6759.1 DSnoT 55.9 48.9 64.1 131.6 114.7 1460.3 3676.6 67845.5 7299247.0 15.3 19.6 146.2 8077.8 6062.6 6948.2 OWL 46.5 42.2 65.7 50.5 62.9 249.4 810.9 19356.5 2302446.2 12.0 15.8 169.1 6811.7 3666.8 4065.4 NeuronAl 48.6 44.9 65.1 55.8 63.6 153.1 52.4 79.3 24807.9 12.6 17.2 123.9 1212.3 551.4 935.0 multiflow Uniform 25.4 28.3 54.7 11.6 13.9 26.0 11.0 13.7 166.9 167.7 341.0 884.3 16.3 19.7 26.8 DSnoT 37.2 42.1 50.2 10.1 12.7 18.2 10.5 13.4 137.6 10.9 14.8 86.0 15.9 19.2 25.3 OWL 23.8 26.7 49.8 10.6 12.9 19.6 10.1 12.7 106.0 84.1 123.0 644.6 16.1 18.5 25.5 NeuronAl 23.7 27.0 42.5 9.9 12.2 17.6 9.8 12.2 72.3 86.1 140.3 562.3 17.0 20.4 28.1 Wanda Uniform 225.8 29.3 48.9 10.7 13.7 24.0 10.8 14.0 122.2 11.3 15.9 101.6 15.2 17.9 23.7 DSnoT 32.2 38.0 50.6 10.4 13.2 20.8 10.8 14.1 109.6 11.4 15.9 96.8 15.8 19.1 24.6 OWL 24.8 28.2 48.6 9.4 11.8 18.5 9.2 11.9 75.1 10.3 14.5 84.5 15.7 17.8 24.5 NeuronAl 25.2 27.2 42.1 9.6 12.0 17.4 9.4 12.0 65.1 9.9 13.8 74.6 16.3 19.1 25.2
Table 7: Perplexity on the 3 Language Modeling datasets computed over 5 different LLMs for four different top-up algorithms (Uniform, DSnoT, OWL, and NeuronAl) on 3 pruning algorithms (Magnitude, multiflow, and Wanda) at 80% sparsity.
Algorithm Top-Up Phi-2.7B LLama-1 7B LLama-2 7B Mistral 7B OPT 6.7B WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB Magnitude Uniform 15320.0 17856.3 31988.6 112885.2 113922.4 139627.3 55800.7 52597.2 89843.5 24815.2 31186.7 7984.1 42902.5 21284.4 22110.8 DSnoT 19977.5 20700.7 18369.5 33994.8 34213.5 72003.2 2358414.2 1802890.0 3015785.0 13297.5 8025.4 5804.0 18112.4 11853.1 14438.1 OWL 6625.1 5597.4 8388.0 169148.8 158523.1 134038.4 26898.0 17892.9 57863.3 9614.4 8503.8 5789.5 33174.7 17756.7 21613.0 NeuronAl 10193.9 9578.7 14427.3 82221.7 85688.7 108410.1 92725.2 67827.9 104616.8 2188.8 1205.3 3226.2 30303.8 20846.6 16807.1 multiflow Uniform 25257.3 12842.4 25929.7 4829.9 2307.4 9806.2 2037.9 1464.8 3878.2 4289.7 2976.3 3812.4 4422.8 2382.7 3276.9 DSnoT 8498.5 3918.3 12343.8 3697.1 2649.8 8263.4 1721.1 1543.0 3443.8 327.0 270.0 752.5 11633.7 9715.6 11764.5 OWL 6255.2 2798.5 13600.8 926.5 563.1 1783.4 544.2 414.3 2819.3 3354.5 2209.7 3555.2 13461.5 11133.9 15080.8 NeuronAl 3550.3 1279.8 5411.4 311.6 188.4 564.8 230.2 269.8 2355.9 983.1 617.5 1467.4 1118.7 591.5 1173.3 Wanda Uniform 20498.5 12422.9 31362.6 5219.5 3969.8 10013.6 4929.7 3117.3 5285.2 330.9 277.7 783.7 4259.6 2345.8 2732.1 DSnoT 15269.0 6855.4 14008.5 3709.7 3076.0 7790.7 5199.4 4443.2 6685.7 346.5 277.3 758.4 7754.1 6157.5 7779.7 OWL 2546.5 1209.5 7061.7 986.5 654.5 2000.4 663.0 486.2 2283.4 206.3 187.8 603.9 13190.1 10605.6 14228.5 NeuronAl 3873.6 2404.5 6675.9 376.1 242.0 566.3 213.9 226.2 993.2 226.3 168.6 698.2 988.2 471.4 988.7
A.2.2 Zero-Shot at 60% and 80% sparsity
In Tables 8-9, we report the results at 60% and 80% sparsity of NeuronAl over Zero-Shot tasks with the five different models tested in the main text. In the first case, our approach turns out to be the best one (when considering the average over the 7 tasks) in 7 out of 10 cases, while for 80% sparsity in 8 out of 10. This further validates the results presented in the main text (for 70% sparsity) since it shows the ability of NeuronAl to provide stable and consistent results over more complicated tasks at different (and high) sparsity levels.
Table 8: Zero-Shot accuracy for 7 tasks, computed over 5 different LLMs for 3 different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on two pruning algorithms (multiflow and Wanda) at 60% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B multiflow 62.09 67.64 63.15 42.04 71.76 39.51 27.2 53.34 w. DSnoT 63.18 66.77 43.49 40.84 66.62 35.41 26.0 48.90 w. OWL 64.62 67.17 60.15 41.78 70.96 37.8 28.8 53.04 w. NeuronAl 62.09 67.40 67.22 42.73 71.04 38.65 27.6 53.82 Wanda 63.54 64.8 69.08 40.16 68.64 34.9 25.4 52.36 w. DSnoT 62.45 64.33 59.17 39.25 64.18 33.53 22.6 49.36 w. OWL 64.62 64.33 64.83 39.80 67.63 34.98 24.2 51.48 w. NeuronAl 63.54 67.09 67.74 40.51 66.54 34.04 24.4 51.98 Llama1 7B multiflow 57.04 62.51 67.19 45.31 59.55 30.97 24.6 49.60 w. DSnoT 49.46 63.06 68.32 44.75 63.80 31.23 26.4 49.57 w. OWL 54.15 63.46 66.54 46.53 60.65 31.14 25.4 49.70 w. NeuronAl 49.10 63.46 68.35 47.89 63.51 32.42 27.8 50.36 Wanda 59.57 62.67 68.81 43.64 62.84 30.38 24.8 50.39 w. DSnoT 51.62 61.64 67.37 43.39 63.89 30.55 25.4 49.12 w. OWL 55.60 64.17 70.61 46.63 62.96 31.74 24.8 50.93 w. NeuronAl 57.76 64.09 70.24 46.59 63.97 31.14 26.2 51.43 Llama2 -7B multiflow 57.04 61.96 64.80 43.39 60.44 29.1 24.6 48.76 w. DSnoT 54.15 63.77 63.91 43.42 66.25 31.83 24.0 49.62 w. OWL 54.87 62.75 65.14 45.20 62.58 29.52 23.8 49.12 w. NeuronAl 52.71 65.90 69.51 46.60 65.87 31.66 24.8 51.01 Wanda 54.15 64.48 65.44 43.85 65.19 30.46 26.2 49.97 w. DSnoT 53.79 64.09 64.83 42.39 63.89 30.03 22.8 48.83 w. OWL 53.79 66.61 66.76 46.63 67.63 32.34 28.0 51.68 w. NeuronAl 53.07 66.06 70.46 46.93 66.67 31.74 26.8 51.68 Mistral-7B multiflow 51.62 49.88 39.17 27.49 29.67 18.77 11.0 32.51 w. DSnoT 54.87 66.61 70.86 45.93 68.27 32.94 24.0 51.93 w. OWL 53.07 50.12 46.33 28.29 32.58 19.28 13.2 34.70 w. NeuronAl 52.71 52.49 52.97 28.77 33.42 18.43 13.2 36.00 Wanda 54.87 66.06 71.13 44.48 67.05 32.00 21.60 51.03 w. DSnoT 54.15 65.59 70.43 44.5 66.88 31.40 20.60 50.51 w. OWL 57.04 67.17 73.85 45.66 67.89 32.59 23.20 52.49 w. NeuronAl 56.32 65.82 70.70 46.71 67.63 33.36 23.20 51.96 OPT-6.7B multiflow 52.71 58.25 62.69 42.24 56.52 25.68 22.20 45.76 w. DSnoT 53.07 58.48 62.57 42.13 57.79 23.98 22.20 45.75 w. OWL 53.07 57.46 63.21 42.98 56.44 25.68 24.20 46.15 w. NeuronAl 52.71 58.64 62.17 41.83 56.31 25.09 22.20 45.56 Wanda 52.71 59.67 62.29 42.80 58.00 25.60 22.6 46.24 w. DSnoT 52.71 58.17 62.11 41.99 57.41 25.43 22.4 45.75 w. OWL 52.71 58.72 62.69 42.14 58.33 25.17 22.6 46.05 w. NeuronAl 52.71 61.25 62.29 42.00 57.49 25.43 23.0 46.31
Table 9: Zero-Shot accuracy for 7 tasks, computed over 5 different LLMs for 3 different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on two pruning algorithms (multiflow and Wanda) at 80% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B multiflow 52.71 50.12 51.04 26.15 27.44 19.37 13.4 34.32 w. DSnoT 50.18 49.49 37.83 26.32 27.82 19.97 13.8 32.2 w. OWL 53.07 51.22 56.76 26.30 30.85 18.52 15.2 35.99 w. NeuronAl 52.71 49.96 50.58 26.49 32.20 19.8 14.4 35.16 Wanda 53.07 49.25 62.17 25.93 27.44 20.99 14.6 36.21 w. DSnoT 53.07 50.99 38.04 26.17 27.10 20.73 13.0 32.73 w. OWL 52.35 51.22 60.21 26.51 29.88 20.05 13.6 36.26 w. NeuronAl 52.71 50.12 62.17 27.57 32.95 21.08 16.0 37.51 Llama-7B multiflow 47.29 50.91 40.03 26.17 26.77 21.16 12.4 32.1 w. DSnoT 46.57 50.43 37.83 26.02 27.06 20.05 12.2 31.45 w. OWL 50.18 50.04 45.47 26.74 27.65 20.48 12.2 33.25 w. NeuronAl 54.87 50.91 55.96 27.62 30.09 18.77 12.8 35.86 Wanda 47.29 49.88 37.83 26.34 26.47 20.99 12.8 31.66 w. DSnoT 46.93 50.36 37.83 26.03 26.56 21.33 13.4 31.78 w. OWL 47.29 49.88 37.83 26.67 27.19 19.54 11.6 31.43 w. NeuronAl 49.82 48.86 40.76 27.41 29.25 18.34 13.0 32.49 Llama2 -7B multiflow 53.43 48.86 37.83 26.35 27.48 21.25 13.2 32.63 w. DSnoT 52.71 48.86 37.86 26.17 26.60 20.39 13.0 32.23 w. OWL 52.71 49.49 37.83 26.62 26.94 19.11 12.6 32.19 w. NeuronAl 52.71 50.36 44.74 27.47 28.62 17.58 14.4 33.70 Wanda 47.65 49.41 37.83 25.82 26.52 20.82 14.6 31.81 w. DSnoT 53.07 47.91 37.86 26.09 27.23 20.73 13.0 32.27 w. OWL 52.71 50.83 37.83 26.52 27.27 19.37 12.8 32.48 w. NeuronAl 52.71 49.49 37.89 27.38 28.58 17.58 14.8 32.63 Mistral-7B multiflow 50.18 48.15 37.80 25.67 26.18 22.70 13.20 31.98 w. DSnoT 52.71 47.36 37.83 26.58 28.03 18.94 13.8 32.18 w. OWL 48.38 49.09 38.44 25.88 25.59 23.29 13.6 32.04 w. NeuronAl 52.71 50.91 37.83 26.35 27.78 20.31 12.8 32.67 Wanda 53.79 48.78 37.83 26.52 27.82 19.8 12.8 32.48 w. DSnoT 52.35 48.30 37.83 26.55 27.44 19.54 13.0 32.14 w. OWL 52.71 47.43 37.83 26.68 27.78 18.52 13.2 32.02 w. NeuronAl 52.71 50.04 38.65 27.52 28.96 19.97 13.4 33.04 OPT-6.7B multiflow 52.71 50.91 37.80 25.87 27.40 19.28 12.40 32.34 w. DSnoT 52.71 50.83 57.31 26.00 25.00 20.22 11.80 34.84 w. OWL 52.71 51.07 37.83 25.74 25.29 20.05 16.00 32.67 w. NeuronAl 52.71 51.70 54.07 26.29 30.64 19.37 11.80 35.23 Wanda 54.15 52.09 41.53 26.47 28.45 18.86 11.80 33.34 w. DSnoT 52.71 51.38 55.32 26.17 27.06 19.37 13.80 35.12 w. OWL 52.71 49.33 37.83 25.84 25.67 20.31 13.00 32.10 w. NeuronAl 52.71 49.25 51.99 26.35 30.39 19.37 12.60 34.67
A.2.3 Zero-Shot at 60%, 70% and 80% sparsity with Magnitude
In Tables 10-12, we report the results of NeuronAl over Zero-Shot tasks using Magnitude pruning as the base pruning algorithm. The results provided by NeuronAl turn out to be the best ones in 9 out of 15 cases. It is also worth noticing that the performance gap between the Magnitude pruning and score-based pruning algorithms (such as Wanda or multiflow) is generally quite high. Hence, NeuronAl can improve the performance of Magnitude (in the standard setting with uniform distribution) to a certain degree since at high sparsity ratios (as the ones we test) the performance of Magnitude has been shown to be poor [21].
Table 10: Zero-Shot accuracy for 7 tasks, computed over 5 different LLMs for 3 different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on Magnitude at 60% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B Magnitude 57.04 62.83 51.38 42.56 66.33 35.41 28.2 49.11 w. DSnoT 54.51 64.09 42.32 41.09 66.25 34.98 26.6 47.12 w. OWL 55.23 62.59 48.81 42.53 67.38 38.48 28.4 49.06 w. NeuronAl 54.15 65.19 47.28 42.36 65.95 36.77 26.8 48.36 Llama-7B Magnitude 51.62 52.64 45.05 39.23 51.05 26.88 20.4 40.98 w. DSnoT 52.35 52.80 46.88 38.3 50.59 26.37 20.6 41.13 w. OWL 52.35 58.41 51.8 42.02 56.31 29.78 23.8 44.92 w. NeuronAl 50.90 56.12 57.43 40.65 55.3 29.78 24.4 44.94 Llama2 -7B Magnitude 51.26 55.8 41.19 36.97 50.17 26.96 16.2 39.79 w. DSnoT 53.79 56.04 42.87 38.3 53.28 27.9 19.8 41.71 w. OWL 51.99 57.3 46.15 42.56 56.65 30.46 19.4 43.5 w. NeuronAl 54.87 59.59 60.4 46.14 59.01 32.85 27.6 48.64 Mistral-7B Magnitude 55.23 62.19 66.36 48.74 67.05 33.19 22.6 50.77 w. DSnoT 55.6 62.35 68.53 48.28 67.51 33.11 23.2 51.23 w. OWL 53.79 64.48 72.17 49.39 68.14 33.87 23.8 52.23 w. NeuronAl 53.79 64.09 71.07 49.83 65.07 35.07 25.0 51.99 OPT-6.7B Magnitude 53.43 50.59 37.86 26.38 26.6 21.42 13.2 32.78 w. DSnoT 52.71 49.25 37.86 26.14 27.27 21.5 13.2 32.56 w. OWL 52.71 50.51 37.83 26.77 30.3 18.52 14.8 33.06 w. NeuronAl 52.71 54.22 39.14 33.3 37.71 24.23 16.8 36.87
Table 11: Zero-Shot accuracy for 7 tasks, computed over 5 different LLMs for 3 different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on Magnitude at 70% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B Magnitude 46.93 53.59 47.22 30.45 47.85 24.57 19.2 38.54 w. DSnoT 46.57 50.91 39.6 30.12 45.54 24.06 16.8 36.23 w. OWL 45.13 52.88 49.2 32.26 51.64 27.56 21.4 40.01 w. NeuronAl 47.65 53.51 54.25 33.4 53.7 29.61 20.6 41.82 Llama-7B Magnitude 53.43 49.96 37.92 27.59 31.73 22.44 16.6 34.24 w. DSnoT 52.71 51.7 37.83 27.71 30.26 22.7 15.4 34.04 w. OWL 53.07 51.38 38.38 33.14 39.31 24.15 16.8 36.6 w. NeuronAl 52.71 54.38 52.91 39.61 46.0 26.62 23.0 42.18 Llama2 -7B Magnitude 51.26 49.96 37.86 25.9 28.45 23.12 13.4 32.85 w. DSnoT 53.79 49.88 37.86 25.42 28.83 20.56 16.6 33.28 w. OWL 53.07 50.28 37.89 26.38 30.77 22.7 15.0 33.73 w. NeuronAl 54.51 55.41 64.86 33.25 42.09 27.47 21.2 42.68 Mistral-7B Magnitude 51.99 50.83 41.13 32.16 42.72 19.54 16.6 36.42 w. DSnoT 53.07 51.62 39.54 31.66 42.51 20.05 16.6 36.44 w. OWL 57.76 56.59 49.17 36.48 45.75 22.01 18.8 40.94 w. NeuronAl 53.79 57.54 62.72 38.68 44.11 26.45 20.6 43.41 OPT-6.7B Magnitude 52.71 49.8 37.83 25.88 26.68 21.33 12.4 32.38 w. DSnoT 52.71 49.96 37.83 25.87 27.19 20.14 13.6 32.47 w. OWL 52.71 50.59 37.83 25.81 25.46 21.25 12.8 32.35 w. NeuronAl 52.71 50.28 37.8 26.27 26.81 20.31 13.0 32.45
Table 12: Zero-Shot accuracy for 7 tasks, computed over 5 different LLMs for 3 different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on Magnitude at 80% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B Magnitude 45.13 50.36 41.19 25.83 29.08 20.9 13.6 32.30 w. DSnoT 46.93 52.33 39.63 25.9 28.32 21.25 13.4 32.54 w. OWL 49.46 50.91 42.35 26.71 35.27 21.67 13.4 34.25 w. NeuronAl 46.57 51.22 44.86 25.78 28.66 21.5 13.0 33.08 Llama-7B Magnitude 46.21 49.96 53.98 25.69 24.83 21.84 13.8 33.76 w. DSnoT 52.35 51.85 38.47 25.52 26.39 21.42 16.0 33.14 w. OWL 48.38 48.93 44.74 25.76 26.35 21.08 15.8 33.01 w. NeuronAl 47.29 48.86 50.34 25.82 26.18 21.25 14.6 33.48 Llama2 -7B Magnitude 52.35 49.57 46.18 25.94 26.14 23.12 16.0 34.19 w. DSnoT 52.71 51.54 37.89 25.46 27.10 22.44 15.4 33.22 w. OWL 53.07 48.70 42.02 25.72 26.60 21.42 14.4 33.13 w. NeuronAl 51.26 49.33 42.14 25.71 26.52 22.70 15.0 33.24 Mistral-7B Magnitude 51.26 50.99 41.16 25.93 27.48 21.84 14.6 33.32 w. DSnoT 52.35 49.72 38.07 26.26 26.43 21.25 14.0 32.58 w. OWL 52.35 50.20 41.04 26.55 27.78 19.97 13.8 33.10 w. NeuronAl 52.71 49.64 43.18 28.11 29.8 21.59 14.6 34.23 OPT-6.7B Magnitude 52.71 49.49 37.83 25.79 26.39 21.25 13.0 32.35 w. DSnoT 52.71 49.57 37.83 25.78 25.63 20.65 12.8 32.14 w. OWL 52.71 49.80 37.83 26.05 26.73 21.16 13.2 32.50 w. NeuronAl 52.71 51.78 37.83 25.79 26.89 21.08 12.8 32.70
A.2.4 NeuronAl on LLama-13B at 70% sparsity
In Table 13, we present the results on LLama-13B at 60% sparsity. The results are in line with the ones at 70% and 80% sparsity presented in the main text since NeuronAl can outperform the competitors in 11 out of 15 cases.
Table 13: Perplexity of LLama-13B on the 3 Language Modeling datasets at 60% sparsity.
Algorithm Top-Up Llama-1 13B LLama-2 13B WikiText2 C4 PTB WikiText2 C4 PTB Magnitude Uniform 34.9 49.1 1413.7 10.1 13.3 457.5 DSnoT 33.6 41.3 604.8 10.1 13.3 376.7 OWL 28.5 36.6 255.0 8.9 11.6 217.2 NeuronAl 24.8 35.1 186.9 9.5 12.3 151.1 multiflow Uniform 8.7 10.9 66.5 16.3 21.0 211.6 DSnoT 8.4 10.8 58.7 8.3 11.3 217.9 OWL 7.9 10.0 48.1 8.5 11.2 120.3 NeuronAl 7.9 10.2 44.7 8.6 11.1 109.9 Wanda Uniform 8.8 11.2 72.1 8.4 11.5 146.0 DSnoT 8.5 11.0 66.4 8.3 11.4 131.3 OWL 7.6 9.8 47.6 7.5 10.2 98.0 NeuronAl 7.6 9.9 46.8 7.6 10.3 90.3
A.2.5 NeuronAl on SparseGPT
In Tables 14-16, we present the results of NeuronAl on SparseGPT on the WikiText2, C4, and PTB datasets, using the block-only setting [^6] Since SparseGPT prunes and updates the weights from columns to rows, the row-wise step of NeuronAl cannot be included: indeed, it would force each row to have a different sparsity ratio, which is in contrast with the nature of SparseGPT..
Table 14: Perplexity on WikiText2 using SparseGPT.
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Uniform 15.8 10.4 10.2 9.4 13.4 OWL 15.8 9.4 9.2 9.1 14.2 NeuronAl 15.7 9.9 9.3 9.1 13.7 70% Uniform 28.9 27.3 27.3 22.0 20.5 OWL 27.7 19.2 20.5 18.6 21.6 NeuronAl 27.3 22.6 20.9 17.8 21.8 80% Uniform 131.0 207.0 122.1 98.4 95.7 OWL 107.5 93.8 84.3 77.2 80.8 NeuronAl 113.5 144.7 88.7 70.8 84.0
Table 15: Perplexity on C4 using SparseGPT.
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Uniform 19.0 12.8 12.9 13.0 15.3 OWL 19.2 11.7 11.6 12.4 15.8 NeuronAl 19.1 12.4 11.7 12.3 15.5 70% Uniform 28.6 28.3 31.5 27.8 22.4 OWL 28.2 21.1 22.8 23.7 22.4 NeuronAl 27.8 23.8 22.5 21.9 22.2 80% Uniform 98.7 136.2 104.8 86.5 72.5 OWL 79.7 68.3 73.4 66.2 65.4 NeuronAl 86.4 104.2 72.4 61.8 65.0
Table 16: Perplexity on PTB using SparseGPT.
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Uniform 28.7 19.5 430.5 73.7 20.3 OWL 29.3 16.9 262.1 70.9 21.0 NeuronAl 28.2 18.2 249.2 67.2 20.6 70% Uniform 50.3 52.6 3780.0 153.2 32.0 OWL 51.0 37.0 1643.4 135.0 32.9 NeuronAl 47.3 40.5 861.6 123.4 32.8 80% Uniform 195.4 295.6 3201.7 316.2 102.3 OWL 141.4 162.3 5556.5 278.8 98.9 NeuronAl 156.7 260.2 3659.8 266.6 105.3
Using SparseGPT, the superiority of NeuronAl is less evident than with other pruning algorithms. Nevertheless, NeuronAl turns out to be the best-performing top-up algorithm in 5 out of 15, 8 out of 15, and 7 out of 15 cases, respectively for WikiText2, C4, and PTB. Interestingly, for lower sparsity, the gap between uniform and non-uniform distribution (both NeuronAl and OWL) is less remarkable than at higher sparsity. We explain these results with the inherent functioning of SparseGPT, which, differently from the other pruning algorithms, includes a weight reconstruction step. However, we can conclude that also in this case, our proposed approach turns out to be effective in many cases at increasing the task performance.
A.3 Block-only and Row-only Ablation
To complement the results presented in the main text on the C4 dataset, Tables 17-18 show the results of the ablation over the block-only and row-only steps also on WikiText2 and PTB. The results are in line with those obtained on C4.
Table 17: Ablation study: perplexity achieved when using the block and row steps alone over the WikiText2 dataset.
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Uniform 25.8 10.7 10.8 11.3 15.2 NeuronAl block-only 25.5 9.4 9.2 9.9 15.2 NeuronAl row-only 25.8 11.2 11.4 11.3 16.5 NeuronAl 25.2 9.6 9.4 9.9 16.3 70% Uniform 227.6 85.1 78.0 60.7 157.5 NeuronAl block-only 108.2 20.3 21.9 29.3 199.1 NeuronAl row-only 200.4 85.3 75.5 60.7 116.8 NeuronAl 89.9 21.5 23.9 29.0 183.9 80% Uniform 20498.5 5219.5 4929.7 330.9 4259.6 NeuronAl block-only 5637.9 458.9 439.9 205.9 1193.0 NeuronAl row-only 11382.1 6715.8 916.8 330.9 4259.6 NeuronAl 3873.6 376.1 213.9 226.3 988.2
Table 18: Ablation study: perplexity achieved when using the block and row steps alone over the PTB dataset.
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Uniform 48.9 24.0 122.2 101.6 23.7 NeuronAl block-only 41.9 16.7 65.6 74.6 23.5 NeuronAl row-only 49.3 22.2 114.5 101.6 26.6 NeuronAl 42.1 17.4 65.1 74.6 25.2 70% Uniform 346.2 157.0 599.3 298.3 209.2 NeuronAl block-only 152.5 41.4 209.7 196.9 250.0 NeuronAl row-only 345.6 148.0 611.8 298.3 170.8 NeuronAl 133.9 43.8 205.3 197.8 194.7 80% Uniform 31362.6 10013.6 5285.2 783.7 2732.1 NeuronAl block-only 6253.3 684.1 1295.2 723.0 1335.6 NeuronAl row-only 18270.0 10008.4 2019.1 783.7 2732.1 NeuronAl 6675.9 566.3 993.2 698.2 988.7
A.4 NeuronAl $\lambda$ Selection
In the main text, we presented an experiment regarding the ability of NeuronAl to pick the most performing $\lambda$ parameters (in the block-only cases) at 70% sparsity. Here we include the same analysis at 60% and 80% sparsity. In Fig.s 9-10, it is clearly visible how NeuronAl still performs correctly over different sparsity ratios. It is also worth noticing that the calibration data always come from the C4 dataset and then the results are transferred to the other unknown datasets.
Figure 9: Perplexity over different values of $\lambda$ at 60 % sparsity. The orange dot indicates the value selected by NeuronAl. The red line corresponds to the perplexity using a uniform distribution, hence $\lambda=0$.
Figure 10: Perplexity over different values of $\lambda$ at 80 % sparsity. The orange dot indicates the value selected by NeuronAl.The red line corresponds to the perplexity using a uniform distribution, hence $\lambda=0$.
A.5 Calibration Data Ablation
In Tables 20-20, we complement the results regarding the seed set of the calibration data at 60% and 80% sparsity. The results are fully in line with the ones presented in the main text. As expected, the standard deviation of the performance increases when increasing the sparsity ratio and at higher sparsity (80%) it turns out to be model dependent.
We also conduct further experiments about the size of the calibration data ($|C_{\lambda}|$), see Fig. 11. Specifically, We compute the perplexity on WikiText2, C4, and PTB using different calibration data of different sizes (1, 2, 4, 8, 16 samples) in order to assess if the performance could be related to it. As clearly visible for almost all models, the performance is basically constant across the different calibration sizes. This does not apply only to OPT models, where for $|C_{\lambda}|=16$ we reach the best performance by a large margin compared to $|C_{\lambda}|=8$ (which is the standard value of $|C_{\lambda}|$ used in all the experiments), especially on WikiText2 and PTB. This, as pointed out in the main text, can better explain the reason why NeuronAl is able to outperform all the top-up competitors in Language Modeling but OPT.
Table 19: Ablation study: perplexity achieved by NeuronAl with different calibration data seeds (0, 16, 46) at 60% sparsity.
Dataset Model
Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B
WikiText2 24.5 $\pm$ 0.6 09.5 $\pm$ 0.1 9.3 $\pm$ 0.0 10.1 $\pm$ 0.1 16.2 $\pm$ 0.1
C4 27.0 $\pm$ 0.2 11.9 $\pm$ 0.1 11.9 $\pm$ 0.0 13.8 $\pm$ 0.1 19.1 $\pm$ 0.1
PTB 42.3 $\pm$ 0.7 17.1 $\pm$ 0.3 65.3 $\pm$ 0.8 74.9 $\pm$ 0.2 25.1 $\pm$ 0.1
Table 20: Ablation study: perplexity achieved by NeuronAl with different calibration data seeds (0, 16, 46) at 80% sparsity.
Dataset Model
Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B
WikiText2 03654.7 $\pm$ 255.1 382.4 $\pm$ 064.7 0247.7 $\pm$ 29.4 216.5 $\pm$ 12.6 1284.9 $\pm$ 482.5
C4 72323.6 $\pm$ 121.7 250.5 $\pm$ 027.1 0265.3 $\pm$ 34.1 171.7 $\pm$ 09.9 0663.5 $\pm$ 316.3
PTB 06014.9 $\pm$ 788.3 624.4 $\pm$ 165.5 1101.9 $\pm$ 94.1 706.1 $\pm$ 06.9 1056.9 $\pm$ 124.5
Figure 11: Perplexity over different values of $|C_{\lambda}|$ (size of the calibration data) when using NeuronAl on the 3 Language Modeling datasets at 70% sparsity.
A.6 NeuronAl $\lambda$ Parameters
Here we show the $\lambda$ parameters selected by our proposed approach for each model, sparsity ratio, and pruning algorithm tested in this work, aiming to facilitate the reproducibility of our results for the community. Please note that such values are the ones used for each combination sparsity-pruning algorithm-model that have been extracted from $C_{\lambda}$ from C4 (using 0 as seed and 8 as size), and then transferred to all the other datasets/tasks. We report the final $\lambda$ values for both the block and row steps in Table 22 for the first 5 models tested in the main text, and in Table 22 for the LLama-13B models.
Table 21: $\lambda$ parameters selected by NeuronAl at sparsity xxx (block |||| row) for each combination sparsity-pruning algorithm-model. Note that for SparseGPT the row step is not possible (see the main text for details).
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Magnitude 0.01 |||| 0.25 0.10 |||| 0.20 0.20 |||| 0.04 0.15 |||| 0.20 0.25 |||| 0.25 Wanda 0.10 |||| 0.25 0.09 |||| 0.25 0.12 |||| 0.20 0.09 |||| 0.00 0.01 |||| 0.15 multiflow 0.08 |||| 0.25 0.12 |||| 0.20 0.12 |||| 0.20 0.08 |||| 0.00 0.01 |||| 0.08 SparseGPT 0.01 |||| 0.25 0.02 |||| 0.25 0.06 |||| 0.25 0.09 |||| 0.02 |||| 0.25 70% Magnitude 0.06 |||| 0.25 0.20 |||| 0.20 0.25 |||| 0.00 0.20 |||| 0.25 0.15 |||| 0.20 Wanda 0.12 |||| 0.25 0.15 |||| 0.20 0.15 |||| 0.25 0.15 |||| 0.25 0.25 |||| 0.20 multiflow 0.15 |||| 0.25 0.15 |||| 0.25 0.12 |||| 0.25 0.15 |||| 0.00 0.25 |||| 0.20 SparseGPT 0.02 |||| 0.25 0.04 |||| 0.25 0.08 |||| 0.25 0.08 |||| 0.25 0.05 |||| 0.25 80% Magnitude 0.01 |||| 0.20 0.03 |||| 0.10 0.01 |||| 0.25 0.15 |||| 0.20 0.20 |||| 0.08 Wanda 0.20 |||| 0.20 0.15 |||| 0.20 0.15 |||| 0.25 0.15 |||| 0.25 0.20 |||| 0.09 multiflow 0.12 |||| 0.25 0.15 |||| 0.15 0.15 |||| 0.25 0.15 |||| 0.20 0.20 |||| 0.05 SparseGPT 0.06 |||| 0.25 0.02 |||| 0.25 0.07 |||| 0.25 0.08 |||| 0.25 0.07 |||| 0.25
Table 22: $\lambda$ parameters selected by NeuronAl on the each combination sparsity-pruning algorithm for Llama-13B (V1 and V2) at sparsity xxx (block |||| row). Sparsity Top-Up Model LLama1 13B LLama2 13B 60% Magnitude 0.10 |||| 0.25 0.12 |||| 0.12 Wanda 0.10 |||| 0.20 0.09 |||| 0.25 multiflow 0.15 |||| 0.20 0.12 |||| 0.20 70% Magnitude 0.12 |||| 0.25 0.25 |||| 0.25 Wanda 0.15 |||| 0.25 0.12 |||| 0.20 multiflow 0.15 |||| 0.15 0.15 |||| 0.25 80% Magnitude 0.25 |||| 0.05 0.25 |||| 0.12 Wanda 0.25 |||| 0.12 0.15 |||| 0.25 multiflow 0.25 |||| 0.10 0.04 |||| 0.25