### v3

# Zeroth-Order Adaptive Neuron Alignment Based Pruning without Re-Training

Abstract
Network pruning focuses on computational techniques that aim to reduce a given model’s computational cost by removing a subset of its parameters while having minimal impact on performance. Throughout the last decade, the most widely used pruning paradigm has been pruning and re-training, which nowadays is inconvenient due to the vast amount of pre-trained models, which are in any case too expensive to re-train. In this paper, we exploit functional information from dense pre-trained models, i.e., their activations, to obtain sparse models that maximize the activations’ alignment w.r.t. their corresponding dense models. Hence, we propose NeuronAl, a top-up algorithm that can be used on top of any given pruning algorithm for LLMs, which modifies the block-wise and row-wise sparsity exploiting information from both the dense model and its sparse version to maximize the neuron alignment among activations. Differently from existing methods, our approach adaptively selects the best hyperparameters for the block-wise and row-wise sparsity ratios w.r.t. the model and the desired sparsity, and requires no re-training. We test our method over 276 cases combining four LLM families, three sparsity ratios, and ten language tasks (three language modeling and seven zero-shot datasets), showing how it consistently outperforms the latest state-of-the-art methods in terms of performance-runtime trade-off. The code is available at https://github.com/eliacunegatti/NeuroAL.
Network Pruning, Large Language Model, Efficiency
[…]
1 Introduction
In recent times, Large Language Models (LLMs) have shown incredible performance over several language tasks (Wei et al., 2022; Min et al., 2023; Chang et al., 2024). However, their performance usually improves with their sizes (i.e., the number of trainable parameters), which in turn is proportional to the computational cost of training and then using such models. One way to reduce this cost is through network pruning, i.e., applying algorithms that remove parameters while minimizing performance degradation. This approach has been extensively studied on Convolutional Neural Networks (CNNs) (Frankle & Carbin, 2019; Lee et al., 2019; Wang et al., 2020; Evci et al., 2020), but nowadays the focus has shifted towards pre-trained models (Touvron et al., 2023a, b; Jiang et al., 2023).
Figure 1: Perplexity vs. Runtime (seconds) trade-off among different top-up algorithms and our proposed NeuronAl based on LLama-7B V1 with a sparsity of 70%, evaluated on WikiText2.
This shift has required a change of paradigm in pruning techniques: in fact, while in CNNs the main paradigm is iterative pruning (with re-training) (Frankle & Carbin, 2019), with pre-trained models (such as LLMs) in most cases it is not possible to fully re-train such models, because (1) training data are often not accessible, and (2) full re-training would be anyway too expensive. This calls for “exploiting” as much as possible the information contained in a pre-trained model to obtain a performant sparse version of it, using weight’s information (Jaiswal et al., 2024), activations (Sun et al., 2023, 2024), or reconstruction error (Frantar & Alistarh, 2023), without the need for re-training. More recently, a new category of pruning algorithms, which we may call top-up algorithms (i.e., methods that can be applied on top of a given pruning algorithm for LLMs), has emerged, aiming at further improving pruning performance. Such approaches can be divided into two categories: those that minimize the reconstruction error (Guo et al., 2024; Xu et al., 2024; Zhang et al., 2024), and those that impose non-uniform sparsity distribution modifying the block-wise sparsity (Yin et al., 2024; Lu et al., 2024; Li et al., 2024). The latter category is extremely effective on CNNs (Frankle et al., 2020; Su et al., 2020), while its application to LLMs has only recently emerged.
Contributions In this paper, we first analyze the major limitations of current top-up algorithms. To do so, we carefully analyze the state-of-the-art top-up methods highlighting their limitations in terms of sensitivity to hyperparameters, the required computational budget, and their block-importance metric. Leveraging this knowledge, we introduce a new top-up method, called NeuronAl. The algorithm consists of a two-step approach that re-distributes the block-wise sparsity, i.e., the sparsity among Transformer blocks, and the row-wise sparsity, i.e., the sparsity for each row of a given layer’s matrix, maximizing a metric which exploits information from both the dense and sparse model, namely the neuron alignment between dense and sparse activations. NeuronAl does not require the user to specify any hyperparameter-tuning, as it automatically selects the most-performing values from a suitable set, hence adapting to the underlying model and the target sparsity. Another advantage is that the neuron alignment only requires the computation of the activations of the dense and sparse models, which reduces the computation budget required, compared to other top-up approaches.
We test our approach on three Language Modeling datasets and seven Zero-Shot tasks over four different LLM families from 7B to 13B parameters, to show its ability to outperform, in the majority of the cases, the most recent state-of-the-art techniques, including OWL, DsNoT (Zhang et al., 2024), and AlphaPruning (Lu et al., 2024) over three different high sparsity values (60%, 70%, and 80%) for a total of 276 test-cases. To assess the robustness of our approach, we also conduct an in-depth sensitivity analysis.
2 Related Work
In this section, we provide a comprehensive discussion about network pruning applied to LLMs. We first introduce structured and unstructured network pruning; then, we focus on the latter, introducing the latest approaches proposed for improving sparse model performance.
Structured Network Pruning. Given a layer’s weight matrix $W\in\mathbb{R}^{n\times m}$ to sparsify, structured pruning removes either entire rows ($n$) or columns ($m$) (see the next section) aiming at speeding up both training and inference time. The first approach that applies structured pruning to LLMs has been proposed in (Ma et al., 2023), and focuses on the dependency of Transformers, i.e., it removes components of the networks while maximizing their original functionality. In (Kurtić et al., 2024), a pruning mechanism has been devised to remove components with the worst balance between loss and runtime. Other structured pruning approaches have been proposed based on combinatorial optimization (Meng et al., 2024), perturbative forward-pass only (Dery et al., 2024), and reduction of the embedding dimension through PCA (Ashkboos et al., 2023). Finally, in (Gromov et al., 2024) it has been found that the last Transformer blocks are redundant, hence they can be completely removed with minor performance drops. The reason behind this phenomenon lies in the similarity between the learnable representation of consecutive blocks, which turns out to increase when the block depth increases. While all these approaches can achieve valuable inference speed-ups, the performance of the resulting sparse models w.r.t. their dense counterparts can be matched only at low sparsity values, such as 20% in (Ma et al., 2023) or 30% in (Ashkboos et al., 2023). This somehow limits the applicability of these methods, since in the case of models with billions of parameters one may need more aggressive pruning strategies to meet stringent hardware requirements.
Unstructured Network Pruning. Differently from structure pruning, unstructured pruning works by removing weights in a scattered (i.e., non-structured) way. While in this scenario the inference speed-up is limited (although techniques for reordering weights are available (Li et al., 2019; Mishra et al., 2021; Zhou et al., 2021)), the performance w.r.t. the dense model can be preserved also at high sparsity ratios (i.e., above 50%), with the performance at lower sparsity being almost always completely preserved. The first approach of this kind has been proposed in (Frantar & Alistarh, 2023), where weight pruning and reconstruction are combined based on the Hessian matrix. Even a simple magnitude-based approach turned out to perform well (Jaiswal et al., 2024), also when integrated with information on the neuron activations (Sun et al., 2023; Farina et al., 2024). These approaches compute a score for each weight and then remove the ones with the lower scores for each layer, with a uniform sparsity across layers.
Top-Up Algorithms To improve the performance of unstructured pruning, several top-up algorithms have been devised. These approaches can be categorized into two distinct groups: methods that minimize the reconstruction error keeping the sparsity uniform for each block, and methods that modify the block-wise sparsity of the model resulting in non-uniform sparsity distribution across blocks.
The first group firstly sparsifies the model using a pruning algorithm and then, either dynamically (Zhang et al., 2024) or by backpropagation (Guo et al., 2024), updates the pruning mask. The second group (to which our method belongs) modifies the block-wise sparsity (obtained by a given pruning algorithm) based either on activations’ outliers (Yin et al., 2024), Empirical Spectral Distance (ESD) (Lu et al., 2024), or allocation functions in a gradient-free manner (Li et al., 2024), while in BESA (Xu et al., 2024) gradient information is used to set layer-wise sparsity using block-wise reconstruction error.
The idea of simply redistributing the layer-wise sparsity is known to be extremely well-performing on Multi-Layer Perceptrons (MLPs) and Convolutional Neural Networks (CNNs). The first approach of this kind, based on the Erdős–Rényi (ER) model, has been proposed in (Mocanu et al., 2018) for MLPs and then adjusted for CNNs in (Evci et al., 2020), while an empirical study about the effect of layer-wise pruning using different sparsity ratios has been done in (Liu et al., 2021). Regarding Transformers (both for vision and text), the state-of-the-art algorithms (Frantar & Alistarh, 2023; Sun et al., 2023) have been devised to set the block-wise sparsity across the Transformer blocks in a uniform way. Later on, OWL (Yin et al., 2024), AlphaPruning (Lu et al., 2024), and DSA (Li et al., 2024) have been proposed to build upon scoring-based pruning algorithms, adjusting the block-wise sparsity in a non-uniform way. These approaches improve the performance of several pruning algorithms, e.g. (Frantar & Alistarh, 2023; Sun et al., 2023), especially at sparsity above 60%. On the same line, BESA (Xu et al., 2024) allocates layer-wise sparsity across each block’s layer using gradient information. Recently, modality-wise sparsity distribution has been investigated in the case of multimodal tasks in (Farina et al., 2024; He & Chen, 2024).
3 Current limitations of top-up algorithms
We discuss below the three main limitations of the state-of-the-art approaches for redistribution of non-uniform sparsity in LLMs, namely 1) their need for hyperparameter tuning, 2) their large runtime, and 3) their block importance metric calculation (hence their block sparsity allocation).
Need for Hyperparameter Tuning We analyze the sensitivity of the hyperparameters used by OWL, namely $\lambda$ and $M$, and by AlphaPruning, namely $\epsilon$. Concerning OWL, the first hyperparameter is used to set how much the sparsity can vary across blocks (i.e., $[s-\lambda,s+\lambda]$) while keeping the overall sparsity fixed as $s$. The second hyperparameter, $M$, defines the outliers’ threshold: namely, for each block, the number of outliers is computed as the number of activations that are $M$ times greater than the block’s activations’ mean. For AlphaPruning, instead, a hyperparameter called $\epsilon$ is used and manually tuned to set two tunable hyperparameters $(s_{1},s_{2})$ that control the sparsity across blocks.
We test the sensitivity of OWL and AlphaPruning to their hyperparameters, using three different sparsity ratios, two LLMs, and Wanda as the underlying pruning algorithm. Fig. 2 displays the perplexity on WikiText2 of the different hyperparameter settings obtained with OWL (first two rows) and AlphaPruning (last row); the gray square corresponds to the best a-posteriori hyperparameter selection. It can be seen that no single hyperparameter value achieves the best performance in all settings, which entails that careful tuning is required for these approaches to be effective.
Figure 2: Perplexity for various hyperparameter settings of OWL ($M$,$\lambda$) and AlphaPruning ($\epsilon$) using Phi-2 and LLama-7b for three sparsity ratios. The gray square corresponds to the hyperparameter values that lead to the best performance.
Large Runtime Another main limitation of some of the current approaches for non-uniform distribution is their computational runtime. This holds mainly BESA (Xu et al., 2024) and DSA (Li et al., 2024). On LLama-7B, the first approach, which relies on gradient information, requires $\sim$5 hours to find the best non-uniform distribution configuration. On the other hand, DSA uses an evolutionary approach to find the best combination of a set of allocation functions, requiring $\sim$12 hours to find the best distribution [^1] The runtime has been taken from the original papers, where authors used A100 GPUs for BESA, and H800 GPUs for DSA.. This high runtime is due to the evaluation of each sparse model obtained by applying all the possible combinations of sparsity allocation functions.
Block Importance Metric Almost all the top-up algorithms select the block-wise sparsity w.r.t. to the block importance given by scoring criteria computed either from the dense model or from the evaluation of the sparse model. In the first case, the block importance is computed on the dense models on a block-wise view using either outliers information, as in OWL, or the ESD, as in AlphaPruning, without focusing on how the selected non-uniform sparsity, once applied, could change the block importance of the successive layers. On the other hand, DSA only uses the perplexity of the sparse model as the evaluation metric. The only exception is BESA, which, similarly to NeuronAl, uses information from both the dense and the sparse model. However, it does so using gradient information, which leads to high computational runtime. Also, DSnoT minimizes the reconstruction error using information from both the dense and sparse models. However, it only updates each layer’s binary mask without changing the sparsity distribution across blocks.
Table 1 summarizes the different properties of the current top-up algorithms and shows how our proposed approach positions w.r.t. the previously proposed approaches.
Table 1: Properties of state-of-the-art top-up algorithms vs. NeuronAl. The runtime is expressed as orders of magnitude computed on LLama-7B. For the metrics, we indicate whether they are computed over the dense ($\mathcal{D}$) and/or sparse ($\mathcal{S}$) models.
Top-Up Hyper. Tuning Runtime (s) Metric DSnoT (Zhang et al., 2024) $10^{0}$ Reconstruction Error ($\mathcal{D},\mathcal{S}$) BESA (Xu et al., 2024) $10^{4}$ Reconstruction Loss ($\mathcal{D},\mathcal{S}$) OWL (Yin et al., 2024) $10^{2}$ Outliers ($\mathcal{D}$) AlphaPruning (Lu et al., 2024) $10^{3}$ ESDs ($\mathcal{D}$) DSA (Li et al., 2024) $10^{4}$ Perplexity ($\mathcal{S}$) NeuronAl (Ours) $10^{2}$ Neuron Alignment ($\mathcal{D},\mathcal{S}$)
4 Methodology
We now describe our proposed NeuronAl algorithm.
Preliminaries Given a dense model $\mathcal{D}$, a pruning algorithm $\mathcal{P}$, a target sparsity $s$, unstructured network pruning generally computes a saliency score $\Psi$ for each weight $w\in\mathcal{D}$ and then binarizes these scores w.r.t. the $\mathtt{top}_{k}$ elements, where $k=1-|\mathcal{D}|\times s$. This allows to obtain a binary mask $\mathcal{M}$ to apply over $\mathcal{D}$, from which the final sparse model can be computed as $\mathcal{S}=\mathcal{D}\odot\mathcal{M}$. Since LLMs are composed of stacked Transformer blocks (each one denoted as $\mathcal{B}_{i}$), i.e., sets of linear layers (each one denoted as $\ell_{i}^{j}$) that implement the self-attention mechanism followed by an MLP, the binarization step is usually done uniformly per each layer $\ell_{i}^{j}$ (Frantar & Alistarh, 2023; Sun et al., 2023) as:
$\mathcal{M}^{\ell_{i}^{j}}=\mathtt{top}_{k^{\ell_{i}^{j}}}(\Psi^{\ell_{i}^{j}},\mathcal{D}^{\ell_{i}^{j}}).$ (1)
Neuron Alignment Our proposed method is based on the idea of combining the concept of neuron alignment, which requires no a priori definition of outliers (hence no $M$ parameter, as in OWL), with that of adaptivity, to remove the need for manually tuning $\lambda$ and $\epsilon$. The method takes as input both $\mathcal{D}$ and its sparse version $\mathcal{S}$ generated by $\mathcal{P}$ with sparsity ratio $s$, and uses a small calibration data $C_{\lambda}$ to make a forward pass on both models, to retrieve the dense and sparse activations, respectively $\mathcal{A}_{\mathcal{D}}$ and $\mathcal{A}_{\mathcal{S}}$. The main idea behind NeuronAl is to maximize the neuron alignment by firstly modifying the vector of sparsity ratios for all blocks ($\mathbf{s}^{\mathcal{B}}$) and then for all rows ($\mathbf{s}^{r}$), where each row corresponds to the layer’s weight matrix $W^{\ell_{i}^{j}}$ (for each layer $\ell_{i}^{j}$ in $\mathcal{B}_{i}$), where $W^{\ell_{i}^{j}}\in\mathbb{R}^{r\times m}$. The main strength of this approach is that it does not require any weight update nor gradient information, but just a block- and row-wise sparsity reallocation and mask update via Eq. (1), using the same scoring criteria of $\mathcal{P}$. It also exploits both the information (i.e., activations) of the original dense model as well as the evaluation of the corresponding sparse model.
However, as tested in the previous observational study, finding the best block/row-wise sparsity requires defining a factor $\lambda$ to control the block/row-wise sparsity difference between consecutive blocks/rows while ensuring the desired global sparsity. While OWL requires $\lambda$ to be set a priori, NeuronAl automatically selects, from a suitable set of values, the best $\lambda$ for each combination of $\mathcal{D}$, $\mathcal{P}$ and $s$, yielding an adaptive top-up method. The only constraint we set is that we use a linear sparsity schedule over $\lambda$ for the block-wise step, demonstrated to be effective in our empirical study in Section 4.3. This choice has been made (1) because we found that the performance improvement obtained with the linear sparsity schedule is more stable, see Table 2, and (2) to align our method to the latest research that shows how the last layers of an LLM have a small influence on the final performance (Gromov et al., 2024).
4.1 Block-Wise Sparsity Ratio
The first step concerns the block-wise redistribution over the whole model. Our method takes as input the dense and sparse models ($\mathcal{D}$ and $\mathcal{S}$), the target sparsity ($s$), the calibration data $C_{\lambda}$, and a set of $\lambda$ parameters ($\lambda^{\text{set}}$). Then, it computes a set of $|\lambda^{\text{set}}|$ vectors of block-wise sparsity values $\mathbf{s}^{\mathcal{B}}_{set}=\{\mathbf{s}^{\mathcal{B}}_{\lambda_{1}},\mathbf{s}^{\mathcal{B}}_{\lambda_{2}},\dots,\mathbf{s}^{\mathcal{B}}_{\lambda_{|\lambda^{\text{set}}|}}\}$, where each element $\mathbf{s}^{\mathcal{B}}_{\lambda_{k}}$ indicates a vector of block-wise sparsity values obtained with a linear schedule in $[s-\lambda_{k},s+\lambda_{k}]$. For each $\mathbf{s}^{\mathcal{B}}_{\lambda_{k}}$, we then forward the calibration data $C_{\lambda}$ through the model, and calculate the corresponding neuron alignment:
$neur_{al}=\sum_{\mathcal{B}_{i}}\sum_{\ell_{i}^{j}}\frac{\Big{\|}\tilde{A}_{D}^{j}-\tilde{A}_{{\mathcal{S}}_{\left(\mathbf{s}^{\mathcal{B}}_{\lambda_{k}}\right)}}^{j}\Big{\|}_{2}}{|\tilde{A}_{D}^{j}|}$ (2)
where $\tilde{A}$ means that the activations are normalized to sum up to one. Then, we select ${\left(\mathbf{s}^{\mathcal{B}}_{set}\right)^{*}}$, i.e., the $\lambda$ parameters per block that minimize Eq. (2). Finally, we update the block-wise sparsity with the selected ${\left(\mathbf{s}^{\mathcal{B}}_{set}\right)^{*}}$ via Eq. (1), thus obtaining a sparsified model $\mathcal{S}_{\mathcal{B}}$.
4.2 Row-Wise Sparsity Ratio
The second step is complementary to the previous one, but in this case, the sparsity is modified w.r.t. the rows of each layer. It is established (Sun et al., 2023) that pruning using the row as a comparison group [^2] Here, comparison group refers to the subset of weights whose scores are compared to decide which weights to prune. achieves better performance w.r.t. using the whole layer since it inherently maximizes the network connectivity (Hoang & Liu, 2023; Cunegatti et al., 2024). Here, we rely on such discovery to strengthen our approach and change the row-wise sparsity based on the neuron alignment of each layer. In this case, for each layer $\ell_{i}^{j}$ (i.e., for each $W^{\ell_{i}^{j}}\in\mathbb{R}^{r\times m}$) we redistribute the sparsity across the $r$ rows. Also in this case the $\lambda$ parameters are critical for deciding how to control the sparsity difference between consecutive rows. We take our sparse model obtained with the block-wise redistribution ($\mathcal{S}_{\mathcal{B}}$) and, for each layer $\ell_{i}^{j}$, we compute different row-wise sparsity values obtaining $\mathbf{s}^{s}_{set}=\{\mathbf{s}^{r}_{\lambda_{1}},\mathbf{s}^{r}_{\lambda_{2}},\dots,\mathbf{s}^{r}_{\lambda_{|\lambda^{\text{set}}|}}\}$, where each $\mathbf{s}^{r}_{\lambda_{k}}$ indicates a vector of row-wise sparsity ranging in $[s-\lambda_{k},s+\lambda_{k}]$, where each element is inversely proportional to the alignment of the corresponding row. In this case, we select in $\mathbf{s}^{s}_{set}$ the row-wise vector ${\left(\mathbf{s}^{r}_{set}\right)^{*}}$ that minimizes Eq. (2), and then apply Eq. (1) to $\mathcal{S}_{\mathcal{B}}$, using each row as comparison group.
The full procedure composed of both the block-wise and row-wise sparsity reallocation is shown in Algorithm 1.
Algorithm 1 Proposed top-up pruning procedure
[…]
Require: $\mathcal{D},\mathcal{P},s,C_{\lambda},\lambda^{\text{set}}$
[…]
$\left(\mathbf{s}^{\mathcal{B}}_{set}\right)^{*}\leftarrow$ NeuronAl ($\mathcal{D},\mathcal{S},\lambda^{\text{set}},C_{\lambda},A_{\mathcal{D}}$) Block step
[…]
$\mathcal{S}_{\mathcal{B}}\leftarrow\mathcal{D}\odot\mathtt{top}_{\left(\mathbf{s}^{\mathcal{B}}_{set}\right)^{*}}(\Psi,\mathcal{D})$
[…]
${\left(\mathbf{s}^{r}_{set}\right)^{*}}\leftarrow\textsc{NeuronAl}(\mathcal{D},\mathcal{S}_{\mathcal{B}},\lambda^{\text{set}},C_{\lambda},A_{\mathcal{D}})$ Row step
[…]
$\mathcal{S}_{\text{final}}\leftarrow\mathcal{D}\odot\mathtt{top}_{\left(\mathbf{s}^{r}_{set}\right)^{*}}(\Psi,\mathcal{D})$
[…]
4.3 Non-Uniform Block-Wise Sparsity Distribution
The sparsity values for each block in $\mathbf{s}^{\mathcal{B}}_{\lambda_{k}}$ have been set such as given the sparsity window $[s-\lambda,s+\lambda]$, these schedules work by redistributing the sparsity across blocks in a monotonically linear way (i.e., the sparsity of block $i$ is always larger than the sparsity of layer $i-1,\forall i>1$). We select this sparsity schedule for two main reasons: (1) as shown below such a straightforward sparsity schedule is already able to achieve similar results w.r.t. state-of-the art approaches, and (2) to align with the latest discovering in the literature of structured pruning where is consistently demonstrate how deeper blocks are redundant and can be removed with marginal performance degradation (Gromov et al., 2024; Men et al., 2024; Kim et al., 2024).
We motivated the choice of a linear schedule by testing three straightforward non-uniform sparsity schedules (namely linear, exponential, and logarithmic), which do not require any block scoring for sparsity allocation.
Table 2 displays the improvement, w.r.t. uniform distribution (averaged across three different Language Modeling datasets, namely WikiText2, C4, PTB), achieved by the three sparsity schedules using Wanda as pruning algorithm with $\lambda=0.08$. The results highlight how non-uniform sparsity schedules, without any block-based scoring, lead to a performance improvement close to OWL’s. Overall, the linear schedule turns out to be the most reliable one since it does not show oscillations in performance across the different sparsity ratios (while this happens for the logarithmic and exponential schedules).
Table 2: Performance improvement w.r.t. uniform distribution averaged across three different datasets (WikiText2, C4, and PTB) using Wanda as pruning algorithm.
Sparsity Model Schedule OWL Exp Log Linear 60% Phi-2.7B +3.4% +2.4% +7.8% +7.7% LLama-1 7B +16.5% +3.4% +15.7% +18.1% 70% Phi-2.7B +45.8% +47.7% +44.9% +52.5% LLama-1 7B +66.8% +28.2% +53.9% +63.5% 80% Phi-2.7B +87.8% +89.3% +55.7% +82.8% LLama-1 7B +81.5% +63.6% -4.4% +68.1% Mean +50.3% +39.1% +28.9% +48.8%
5 Experiments
We apply our proposed NeuronAl to different state-of-the-art pruning algorithms tailored for LLMs. Specifically, we test how it compares in terms of performance over Language Modeling datasets and Zero-Shot tasks w.r.t. the most recent top-up algorithms for pruning. We also perform scalability and sensitivity analyses to show the effectiveness of our NeuronAl.
5.1 Experimental Setup
Language Modeling Datasets To measure the models’ perplexity on Language Modeling datasets, we use the following three datasets: (1) WikiText2 (Merity et al., 2017), (2) Colossal Clean Common Crawl (C4) (Raffel et al., 2020), and (3) Penn Treebank (PTB).
Zero-Shot Tasks To assess more thoroughly how the different pruning algorithms affect the models’ capabilities, we employ the following 7 datasets: (1) Recognizing Textual Entailment (RTE) (Dagan et al., 2006; Bar Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009) , (2) WinoGrande (Sakaguchi et al., 2021), (3) BoolQ (Clark et al., 2019), (4) HellaSwag (Zellers et al., 2019), (5) ARC-e (Clark et al., 2018), (6) ARC-c (Clark et al., 2018), (7) OBQA (Mihaylov et al., 2018)
Table 3: Perplexity on the three Language Modeling datasets computed over five different LLMs for four different top-up algorithms (Uniform, DSnoT, OWL, and NeuronAl) on three pruning algorithms (Magnitude, multiflow, and Wanda) at 70% sparsity.
Algorithm Top-Up Phi-2.7B LLama-1 7B LLama-2 7B Mistral 7B OPT 6.7B WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB Magnitude Uniform 764.6 384.4 983.9 2.53e4 2.25e4 3.26e4 1.42e5 1.02e4 2.02e6 221.9 232.9 748.7 1.00e4 5.39e3 6.54e3 DSnoT 539.0 258.0 656.2 1.02e7 2.77e6 4.99e7 1.31e8 2.90e7 2.25e8 192.7 189.9 566.2 6.16e3 3.93e3 4.36e3 OWL 419.6 242.7 358.5 1.20e4 6.58e3 5.39e4 3.39e5 1.24e4 3.28e6 111.7 124.2 545.5 1.57e4 8.48e3 9.67e3 AlphaPruning 2.52e4 1.60e4 2.34e4 424.9 391.5 5.08e4 3.37e3 3.60e3 1.73e5 91.3 106.5 717.1 1.22e4 7.22e3 7.51e3 NeuronAl 281.7 180.9 321.1 231.8 219.9 4.46e3 155.8 264.8 2.61e3 46.5 43.1 612.8 2.11e4 1.07e4 1.09e4 multiflow Uniform 388.4 298.8 610.8 80.9 71.9 172.4 60.0 58.8 1.26e3 9.37e2 6.56e2 2.06e3 9.44e2 1.25e3 843.1 DSnoT 325.5 261.9 328.8 67.6 65.0 114.7 66.6 75.8 6.89e2 57.4 63.3 2.65e2 241.8 153.3 263.9 OWL 197.9 141.3 293.9 25.1 25.8 78.9 29.2 31.0 5.47e2 329.0 7.64e2 1.72e3 240.9 495.6 337.8 AlphaPruning 1.22e5 8.99e4 9.52e4 32.2 35.2 103.8 31.3 34.0 287.3 230.8 292.8 1.72e3 133.8 63.7 153.9 NeuronAl 105.4 87.1 179.5 20.7 21.2 46.2 22.1 23.9 265.5 202.5 334.7 1.41e3 209.7 83.7 202.1 Wanda Uniform 227.6 182.7 346.2 85.1 86.2 157.0 78.0 81.0 599.3 60.7 73.6 298.3 157.5 260.1 209.2 DSnoT 221.9 172.6 257.6 72.9 76.0 121.0 76.1 85.7 491.8 81.3 79.9 304.8 191.4 173.3 182.6 OWL 132.7 116.2 183.7 24.6 27.3 61.2 30.5 36.6 333.7 41.0 51.8 253.5 54.4 69.7 100.7 AlphaPruning 4.22e4 3.05e4 2.23e4 26.9 31.1 77.4 32.0 37.7 273.8 39.4 49.8 286.8 93.8 53.7 120.9 NeuronAl 88.3 77.7 129.5 21.5 23.2 44.2 24.0 27.4 207.0 28.8 33.7 232.0 172.6 84.0 182.7
Models and Sparsity Since one of the distinctive features of NeuronAl is its adaptability w.r.t. sparsity and models, we test four different LLM families, namely LLama-7B (both V1 and V2) (Touvron et al., 2023a, b), Phi-2, Mistral-7B (Jiang et al., 2023), and OPT-6.7B (Zhang et al., 2022). To scale up the model size, we also test LLama-13 (both V1 and V2). In the paper, we mainly present results at 70% sparsity, for fair comparisons with (Yin et al., 2024; Lu et al., 2024). To assess the generalization to different sparsity ratios, we also include 60% and 80% sparsity in our experiments, see the Appendix.
Baselines As pruning algorithms, we test Magnitude, multiflow (Farina et al., 2024), and Wanda (Sun et al., 2023). All are tested with four different top-up algorithms (besides ours): (1) Uniform distribution, (2) DsnoT (Zhang et al., 2024) (dynamic training-free uniform distribution with mask update), (3) OWL (Yin et al., 2024) (block-wise training-free non-uniform distribution based on outliers scores) and (4) AlphaPruning (Lu et al., 2024) (block-wise training-free non-uniform distribution based on ESD) [^3] BESA and DSA are not included in these experiments due their large runtime. Testing them on all combinations of sparsities and pruning algorithms is unfeasible with our GPU resources.. For OWL, we set the hyperparameters to the values that are used mostly in the original paper, hence $M=5$ and $\lambda=0.08$; we do the same for AlphaPruning, setting $\epsilon=0.3$. All these baselines are tested considering each row as a comparison group: in other words, for a given layer, the sparsity $s$ is uniform for each row of each matrix rather than uniform across matrices. This is done for two main reasons: 1) as mentioned earlier, it is established that row-wise pruning on LLMs leads to better performance w.r.t. layer-wise pruning (Sun et al., 2023), and 2) since our approach relies on a row-wise step, for fairness we also use each row (rather than layer) as a comparison group on all the other methods, to appreciate the benefit of our approach [^4] We also test our approach on SparseGPT (Frantar & Alistarh, 2023), using in this case only the block-wise approach, since SparseGPT relies on a weight reconstruction mechanism that firstly prunes columns and then updates the rows of the pruned cells, which makes it unfeasible to apply our row-wise approach. The results can be found in the Appendix 18-20.. For all the pruning algorithms that use calibration data (i.e., multiflow, Wanda, and SparseGPT), we use 128 samples from the C4 dataset, as in (Frantar & Alistarh, 2023; Sun et al., 2023; Yin et al., 2024).
NeuronAl Setup Our method takes as input an LLM model, a target sparsity, a scoring-based pruning algorithm, and two sets of $\lambda$ parameters (for the block- and the row-wise steps, respectively). In the experiments, we set $\lambda^{\text{set}}$ = [0.01, 0.02, 0.03, 0.05, 0.06, 0.07, 0.08, 0.09, 0.1, 0.12, 0.15, 0.20,0.25] for the block-wise step, while for the row-wise step, we also added $0.0$ (in case of no performance improvement). Each value in $\lambda^{\text{set}}$ has been evaluated (as described in Algorithm 1) over a calibration data $C_{\lambda}$. Since all the base pruning algorithms require a single forward over $C$ (with $C$ containing 128 sequences of 2048 tokens each), while OWL requires a second forward always over $C$, to make the computational runtime similar we set $C_{\lambda}=8$ (the closest power of 2 w.r.t. $|C|/|\lambda^{\text{set}}|$). In essence, NeuronAl only requires two forward steps more than the base pruning algorithms, and one forward step more than OWL [^5] For both $C$ and $C_{\lambda}$, we use the same seed (00) for the calibration set, i.e., $C_{\lambda}$ contains the first 8 elements of $C$.. All the experiments have been run on NVIDIA A100 GPUs, both with 40 and 80 GB.
5.2 Experimental Results
Concerning the Language Modeling datasets, the numerical results in terms of perplexity computed over the three Language Modeling datasets at 70% sparsity are shown in Table 3. It can be seen how NeuronAl is able in almost all cases to outperform all the other baselines by a large margin. In no case NeuronAl performs worse w.r.t. the uniform distribution. The only model on which NeuronAl is not the best top-up algorithm for all pruning algorithms is OPT. In all other cases, NeuronAl outperforms OWL and AlphaPruning for all models and pruning algorithms. The results at 60% and 80% sparsity shown in Tables 9-10 in the Appendix confirm this trend.
As for the Zero-Shot tasks, the numerical results are shown in Table 4-5666Magnitude is omitted from these tables, due to space constraints. Results are available in Table 14-16 in the Appendix.. We display only the mean across the seven Zero-Shot tasks, while the results for each task are available in 11-13 in the Appendix. Again, NeuronAl turns out to outperform in the majority of cases all the baselines. In 20 cases out of 30 (w.r.t. the mean accuracy across all tasks), NeuronAl is the one that reaches the best performance and in 5 cases the second best.
Table 4: Accuracy on the seven Zero-Shot Tasks using Wanda as pruning algorithm.
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Uniform 52.36 50.39 49.97 51.03 46.24 DSnoT 49.36 49.12 48.83 50.51 45.75 OWL 51.48 50.93 51.68 52.49 46.05 AlphaPruning 43.14 51.08 51.21 49.87 44.50 NeuronAl 52.04 51.41 51.99 52.09 46.10 70% Uniform 39.89 36.90 34.37 36.85 36.31 DSnoT 38.76 36.26 34.09 36.53 36.35 OWL 41.20 43.31 40.57 38.77 38.77 AlphaPruning 35.02 44.08 42.03 39.05 39.53 NeuronAl 41.87 44.57 43.10 41.56 38.86 80% Uniform 36.21 31.66 31.81 32.48 33.34 DSnoT 32.73 31.78 32.27 32.14 35.12 OWL 36.26 31.43 32.48 32.02 32.10 AlphaPruning 30.74 35.34 32.09 32.29 32.16 NeuronAl 37.36 36.31 32.74 33.08 33.05
Table 5: Accuracy on the seven Zero-Shot Tasks using multiflow as pruning algorithm.
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Uniform 53.34 49.60 48.76 32.51 45.76 DSnoT 48.90 49.57 49.62 51.93 45.75 OWL 53.04 49.70 49.12 34.70 46.15 AlphaPruning 40.25 49.13 48.95 34.97 45.01 NeuronAl 53.77 50.57 51.08 34.07 45.51 70% Uniform 38.40 38.51 36.98 32.76 33.91 DSnoT 38.76 37.51 36.25 37.83 36.33 OWL 42.46 43.23 40.60 32.63 35.69 AlphaPruning 31.81 44.07 41.50 31.95 37.83 NeuronAl 43.18 45.93 42.87 32.63 39.34 80% Uniform 34.32 32.10 32.63 31.98 32.34 DSnoT 32.20 31.45 32.23 32.18 34.84 OWL 35.99 33.25 32.19 32.04 32.67 AlphaPruning 30.72 36.42 32.30 32.77 32.65 NeuronAl 34.86 37.61 32.99 32.32 36.63
Table 6: Perplexity of LLama-13B on the three Language Modeling datasets at 70% (top) and 80% sparsity (bottom).
Algorithm Top-Up Llama-1 13B LLama-2 13B WikiText2 C4 PTB WikiText2 C4 PTB multiflow Uniform 49.4 45.3 277.8 144.3 112.4 623.2 DSnoT 46.2 48.9 240.4 45.8 54.2 611.5 OWL 16.6 17.7 132.2 54.0 56.2 426.6 AlphaPruning 13.8 16.1 132.6 26.9 32.6 337.9 NeuronAl 13.8 15.6 101.7 20.1 23.1 318.9 Wanda Uniform 54.4 55.3 309.2 45.7 56.2 571.0 DSnoT 47.8 54.2 248.6 46.6 57.7 555.5 OWL 16.3 18.9 147.6 18.0 21.8 315.1 AlphaPruning 14.6 17.3 126.7 15.2 18.8 271.1 NeuronAl 14.3 16.6 97.9 16.5 19.3 237.6 multiflow Uniform 3.71e3 1.70e3 3.59e3 4.48e3 2.41e3 5.21e3 DSnoT 5.37e3 2.86e3 6.29e3 1.94e3 1.67e3 5.28e3 OWL 813.8 375.7 2.14e3 1.80e3 1.01e3 4.39e3 AlphaPruning 210.7 147.8 1.19e3 458.1 279.7 1.93e3 NeuronAl 126.8 123.7 901.6 1.60e3 864.1 2.96e3 Wanda Uniform 3.48e3 1.96e3 3.57e3 1.12e3 870.5 5.55e3 DSnoT 4.37e4 2.44e4 3.22e4 4.44e3 3.96e3 4.09e3 OWL 761.6 368.1 1.93e3 248.0 204.2 2.03e3 AlphaPruning 209.6 148.6 973.3 165.1 158.2 1.53e3 NeuronAl 156.4 132.4 1.47e3 185.7 143.1 1.25e3
Scalability Study To assess if the NeuronAl performance scales to bigger models, we apply it to LLama-13B (both V1 and V2) on the Language Modeling datasets. The results available in Table 6 (top) show how our approach is even more effective with larger models. For the 70% sparsity ratio, NeuronAl turns out to reach the best performance in 10 out of 12 cases, being the second best otherwise, for all pruning algorithms. The gap in performance w.r.t. the other baselines (in particular uniform distribution) is even more pronounced than in the smaller models tested previously: in fact, NeuronAl performs best in the majority of the cases (being second best otherwise), with AlphaPruning being close in terms of performance.
Figure 3: Perplexity over different values of $\lambda$ at 70% sparsity. The orange dot indicates the value selected by NeuronAl.
Runtime vs. Perplexity NeuronAl provides a good trade-off between performance and runtime. In Table 7, we show for all baselines (here we also include BESA and DSA), the runtime in seconds required to obtain the non-uniform sparsity distribution for the given model (in this case LLama-7B V1) as well as the performance computed as the perplexity over WikiText2. The results confirm how NeuronAl can achieve, in 3 out of 4 cases, the best results in terms of perplexity while maintaining a low computational budget. In terms of runtime, the only comparable methods are DSnoT and OWL, compared to which however NeuronAl achieves better performance. On the other hand, DSA is the closest in terms of perplexity to NeuronAl, while requiring four orders of magnitude more time to obtain the best sparsity distribution. Overall, the performance-runtime trade-off of NeuronAl improves when increasing the sparsity ratio.
Table 7: Runtime (seconds) vs. perplexity trade-off comparison among different top-up algorithms over LLama-7B pruned at different sparsity ratios using Wanda.
Metric Top-up pruning algorithms Uniform DsNoT OWL BESA DSA AlphaPruning NeuronAl Runtime - 4.5s 73.3s $\sim$ 1.8 ×$\times 10^{4}$ s $\sim$ 4.3 ×$\times 10^{4}$ s 1479.4s 237.1s Perplexity @ 65% 20.9 19.1 13.1 18.5 12.6 14.0 12.8 Perplexity @ 70% 85.1 72.9 24.6 42.6 22.6 26.9 20.7 Perplexity @ 75% 927.4 646.7 152.5 257.9 103.3 110.2 61.2 Perplexity @ 80% 5.22e3 3.71e3 986.5 2.21e3 736.81 768.4 302.8
5.3 Further Analyses
NeuronAl $\lambda$ Selection In this section, we report an analysis of the ability of NeuronAl to pick the best $\lambda$ parameters (i.e., the parameters for which the performance is the best one, hence the lowest value if computed over perplexity). To do this, we evaluate NeuronAl for all the $\lambda$ parameters (in the block-only scenario to simplify the visualization of results) over the three Language Modeling datasets. Fig. 3 reports the perplexity at 70% sparsity across different values of $\lambda$ (black dots connected by solid lines), while the dots highlighted in orange indicate the perplexity achieved with the $\lambda$ value selected by NeuronAl. These results highlight how NeuronAl, in the majority of the cases, can pick the best value of $\lambda$ both with data knowledge, as in the C4 dataset (from which the calibration data is sampled), as well as on unseen datasets such as WikiText2 and PTB. Fig.s LABEL:fig:lambda06-LABEL:fig:lambda08 in the Appendix show the results at 60%-80% sparsity.
Sensitivity to Calibration Data Since NeuronAl works uses a calibration set, we test its sensitivity on the calibration settings in terms of seed, data source, and size. The results in Table 8 show the mean and standard deviation of perplexity at 70% sparsity over the three Language Modeling datasets for different seeds and when using different datasets to extract the calibration data. Furthermore, the results in Fig. 4 report the perplexity at 70% sparsity over the three Language Modeling datasets when changing the number of calibration samples (using C4 as calibration source and 0 as seed).
Overall, it can be seen that our method is fairly robust w.r.t. the calibration source, seeds, and number of samples. The only exception is the OPT model, where the sensitivity to the calibration data turns out to be higher. This sensitivity analysis reveals a possible explanation for the results presented in Section 5.2, where our proposed approach consistently outperformed the baselines over all models and tasks, except for the OPT family.
Table 8: Perplexity (avg. $\pm$ std. dev.) achieved when using NeuronAl with different calibration data seeds (top), and when using different datasets as the source for the calibration data (bottom) on the three Language Modeling datasets at 70% sparsity.
[…]
Figure 4: Perplexity over different values of $|C_{\lambda}|$ (size of the calibration data) when using NeuronAl on the three Language Modeling datasets at 70% sparsity.
6 Conclusion and Limitations
In this paper, we proposed NeuronAl, a new approach to prune LLMs based on the neuron alignment between sparse and dense activations. The main novelty of our approach is that it exploits information from both the dense and the sparse models while also being adaptive since it is designed to automatically select the best hyperparameters for a given model, pruning algorithm, and target sparsity. Throughout extensive experiments, we showed how our approach outperforms, in most cases, the latest state-of-the-art methods both on Language Modeling datasets and Zero-Shot tasks, with different LLM families and sparsity ratios, while requiring minimal time overhead w.r.t. the base pruning algorithm applied. We also included an extensive sensitivity analysis to show the robustness of our approach to the calibration data, and its capability to select $\lambda$, as well as a runtime comparison with the most recent competitors.
The present version of NeuronAl has two main limitations. (1) With NeuronAl, it is not possible to make use of optimized semi-structured sparsity inference implementations (e.g, the NVIDIA N:M sparsity (Pool, 2020)): for a given sparsity, NeuronAl produces customized sparsity constraints for each layer in a block, and then for each row of each layer. Therefore, these implementations cannot be employed as they often require continuity in the sparsity of matrices. (2) Secondly, it requires the definition of $\lambda^{\text{set}}$. However, it is worth noticing that $|\lambda^{\text{set}}|$ as well as $|\mathcal{C}_{\lambda}|$ affect the computational cost of the forward step.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 Additional experiments
Here we include the results of the experiments that, due to space limits, we could not include in the main text. Specifically, we report: the results of NeuronAl on the Language Modeling tasks at 60% and 80% sparsity and the full results on the Zero-Shot tasks at 60%, 70%, and 80% sparsity; the results for the Zero-Shot tasks with Magnitude pruning; the results over LLama-13B at 70% sparsity; the results of NeuronAl (block-only) applied to SparseGPT (Frantar & Alistarh, 2023).
A.1.1 Language Modeling at 60% and 80% sparsity
In Table 9-10, we report the results of NeuronAl over the 3 Language Modeling datasets (WikiText2, C4, and PTB) with the five different models considered in the main text, for 60% and 80% sparsity. In the first case, our approach turns out to be the best one in 23 out of 45 cases, while for 80% sparsity in 20 out of 45, while is second best in 15 cases. It is interesting to notice how at medium sparsity (60%) all the top-up algorithms, including ours, provide similar results, while the improvement provided by NeuronAl at 80% (w.r.t. the top-up competitors) in some cases reaches a factor of 2-3x (e.g., with Llama-1 7B for multiflow and Wanda).
Table 9: Perplexity on the three Language Modeling datasets computed over five different LLMs for four different top-up algorithms (Uniform, DSnoT, OWL, and NeuronAl) on three pruning algorithms (Magnitude, multiflow, and Wanda) at 60% sparsity.
Algorithm Top-Up Phi-2.7B LLama-1 7B LLama-2 7B Mistral 7B OPT 6.7B WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB Magnitude Uniform 51.3 45.9 66.9 152.4 159.8 3.02e3 6.89e3 4.27e4 1.71e6 19.6 24.4 189.3 9.49e3 6.20e3 6.76e3 DSnoT 55.9 48.9 64.1 131.6 114.7 1.46e3 3.68e3 6.78e4 7.30e6 15.3 19.6 146.2 8.08e3 6.06e3 6.95e3 OWL 46.5 42.2 65.7 50.5 62.9 249.4 810.9 1.94e4 2.30e6 12.0 15.8 169.1 6.81e3 3.67e3 4.07e3 AlphaPruning 268.7 254.2 1.13e3 61.6 82.2 153.6 37.1 49.4 5.43e5 13.1 18.2 185.4 5.07e3 2.62e3 3.03e3 NeuronAl 48.5 44.8 65.1 55.8 63.6 153.1 52.4 79.3 2.48e4 12.5 17.0 139.0 1.20e3 547.6 932.6 multiflow Uniform 25.4 28.3 54.7 11.6 13.9 26.0 11.0 13.7 166.9 167.7 341.0 884.3 16.3 19.7 26.8 DSnoT 37.2 42.1 50.2 10.1 12.7 18.2 10.5 13.4 137.6 10.9 14.8 86.0 15.9 19.2 25.3 OWL 23.8 26.7 49.8 10.6 12.9 19.6 10.1 12.7 106.0 84.1 123.0 644.6 16.1 18.5 25.5 AlphaPruning 434.1 401.2 1.82e3 11.6 14.1 19.9 10.9 13.8 83.4 89.1 116.4 454.6 30.9 25.4 45.3 NeuronAl 23.6 27.0 42.0 9.9 12.2 17.7 9.7 12.1 72.2 112.4 168.0 806.9 16.6 19.9 27.3 Wanda Uniform 225.8 29.3 48.9 10.7 13.7 24.0 10.8 14.0 122.2 11.3 15.9 101.6 15.2 17.9 23.7 DSnoT 32.2 38.0 50.6 10.4 13.2 20.8 10.8 14.1 109.6 11.4 15.9 96.8 15.8 19.1 24.6 OWL 24.8 28.2 48.6 9.4 11.8 18.5 9.2 11.9 75.1 10.3 14.5 84.5 15.7 17.8 24.5 AlphaPruning 165.3 166.5 669.7 10.1 12.7 17.7 9.8 12.6 69.3 10.8 15.3 93.5 27.9 24.3 41.2 NeuronAl 25.3 27.1 41.8 9.6 12.0 17.4 9.4 12.0 64.3 9.9 13.8 81.9 16.3 19.1 25.2
Table 10: Perplexity on the three Language Modeling datasets computed over five different LLMs for four different top-up algorithms (Uniform, DSnoT, OWL, and NeuronAl) on three pruning algorithms (Magnitude, multiflow, and Wanda) at 80% sparsity.
Algorithm Top-Up Phi-2.7B LLama-1 7B LLama-2 7B Mistral 7B OPT 6.7B WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB WikiText2 C4 PTB Magnitude Uniform 1.53e4 1.79e4 3.20e4 1.13e5 1.14e5 1.40e5 5.58e4 5.26e4 8.98e4 2.48e4 3.12e4 7.98e3 4.29e4 2.13e4 2.21e4 DSnoT 1.99e4 2.07e4 1.84e4 3.40e4 3.42e4 7.20e4 2.36e6 1.80e6 3.02e6 1.33e4 8.03e3 5.80e3 1.81e4 1.19e4 1.44e4 OWL 6.63e3 5.60e3 8.39e3 1.69e5 1.59e5 1.34e5 2.69e4 1.79e4 5.79e4 9.61e3 8.50e3 5.79e3 3.32e4 1.78e4 2.16e4 AlphaPruning 8.54e4 9.37e4 7.83e4 5.78e4 4.54e4 9.22e4 4.83e4 2.46e4 5.09e4 3.03e3 2.74e3 3.30e3 2.52e4 1.31e4 1.32e4 NeuronAl 9.34e3 8.36e3 1.14e4 8.23e4 8.57e4 1.08e5 3.64e4 3.15e4 2.99e4 1.42e3 7.53e2 4.45e3 3.42e4 2.27e4 1.84e4 multiflow Uniform 2.53e4 1.28e4 2.59e4 4.83e3 2.31e3 9.81e3 2.04e3 1.46e3 3.88e3 4.29e3 2.98e3 3.81e3 4.42e3 2.38e3 3.28e3 DSnoT 8.50e3 3.92e3 1.23e4 3.70e3 2.65e3 8.26e3 1.72e3 1.54e3 3.44e3 327.0 270.0 752.5 1.16e4 9.72e3 1.18e4 OWL 255.2 2.80e3 1.36e4 926.5 563.1 1.78e3 544.2 414.3 2.82e3 3.35e3 2.21e3 3.56e3 1.35e4 1.11e4 1.51e4 AlphaPruning 2.12e4 1.27e4 2.04e4 934.0 841.6 1.55e3 899.1 670.9 2.60e3 3.46e3 3.69e3 6.05e3 4.21e3 2.92e3 3.48e3 NeuronAl 2.34e3 992.4 4.22e3 259.8 209.2 613.8 378.5 456.8 2.09e3 1.02e3 719.3 1.56e3 1.29e3 721.9 1.35e3 Wanda Uniform 2.05e4 1.24e4 3.14e4 5.22e3 3.97e3 1.00e4 4.93e3 3.12e3 5.29e3 330.9 277.7 783.7 4.26e3 2.35e3 2.73e3 DSnoT 1.53e4 6.86e3 1.40e4 3.71e3 3.08e3 7.79e3 5.20e3 4.44e3 6.69e3 346.5 277.3 758.4 7.75e3 6.16e3 7.78e3 OWL 2.55e3 1.21e3 7.06e3 986.5 654.5 2.00e3 663.0 486.2 2.28e3 206.3 187.8 603.9 1.32e4 1.06e4 1.42e4 AlphaPruning 4.31e4 3.33e4 2.66e4 768.4 654.9 1.29e3 982.1 670.0 2.18e3 204.3 182.6 774.2 5.61e3 4.97e3 5.21e3 NeuronAl 2.50e3 1.59e3 4.03e3 302.8 272.2 783.8 557.2 660.2 2404.1 249.5 177.6 783.9 1.04e3 632.8 1.13e3
A.1.2 Zero-Shot at 60%, 70%, and 80% sparsity
In Tables 11-13, we report the results at 60%, 70%, and 80% sparsity of NeuronAl over Zero-Shot tasks with the five different models tested in the main text. Here we report the detailed results for each task, while in the main paper, we only reported the average results across the seven tasks.
Table 11: Accuracy on the seven Zero-Shot Tasks, computed over five different LLMs for three different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on two pruning algorithms (multiflow and Wanda) at 60% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to the uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B multiflow 62.09 67.64 63.15 42.04 71.76 39.51 27.2 53.34 w. DSnoT 63.18 66.77 43.49 40.84 66.62 35.41 26.0 48.90 w. OWL 64.62 67.17 60.15 41.78 70.96 37.8 28.8 53.04 AlphaPruning 62.45 52.96 50.83 29.03 48.27 22.18 16.0 40.25 w. NeuronAl 62.82 67.4 65.87 42.96 70.88 38.65 27.8 53.77 Wanda 63.54 64.8 69.08 40.16 68.64 34.9 25.4 52.36 w. DSnoT 62.45 64.33 59.17 39.25 64.18 33.53 22.6 49.36 w. OWL 64.62 64.33 64.83 39.80 67.63 34.98 24.2 51.48 AlphaPruning 64.26 55.33 62.39 31.07 47.9 23.81 17.2 43.14 w. NeuronAl 64.26 66.38 67.55 40.63 66.5 34.39 24.6 52.04 Llama1 7B multiflow 57.04 62.51 67.19 45.31 59.55 30.97 24.6 49.60 w. DSnoT 49.46 63.06 68.32 44.75 63.80 31.23 26.4 49.57 w. OWL 54.15 63.46 66.54 46.53 60.65 31.14 25.4 49.70 AlphaPruning 50.9 65.43 67.65 46.11 58.08 31.57 24.2 49.13 w. NeuronAl 50.9 63.54 68.35 47.91 63.89 32.17 27.2 50.57 Wanda 59.57 62.67 68.81 43.64 62.84 30.38 24.8 50.39 w. DSnoT 51.62 61.64 67.37 43.39 63.89 30.55 25.4 49.12 w. OWL 55.60 64.17 70.61 46.63 62.96 31.74 24.8 50.93 AlphaPruning 59.57 65.27 68.81 46.09 60.4 32.85 24.6 51.08 w. NeuronAl 58.48 63.61 70.55 46.53 63.8 30.89 26.0 51.41 Llama2 -7B multiflow 57.04 61.96 64.80 43.39 60.44 29.1 24.6 48.76 w. DSnoT 54.15 63.77 63.91 43.42 66.25 31.83 24.0 49.62 w. OWL 54.87 62.75 65.14 45.20 62.58 29.52 23.8 49.12 AlphaPruning 52.71 64.96 64.86 44.63 61.32 30.8 23.4 48.95 w. NeuronAl 53.07 65.27 69.27 46.85 66.62 31.31 25.2 51.08 Wanda 54.15 64.48 65.44 43.85 65.19 30.46 26.2 49.97 w. DSnoT 53.79 64.09 64.83 42.39 63.89 30.03 22.8 48.83 w. OWL 53.79 66.61 66.76 46.63 67.63 32.34 28.0 51.68 AlphaPruning 54.15 67.4 66.67 46.17 64.98 33.28 25.8 51.21 w. NeuronAl 52.71 66.77 71.99 46.85 66.33 32.08 27.2 51.99 Mistral-7B multiflow 51.62 49.88 39.17 27.49 29.67 18.77 11.0 32.51 w. DSnoT 54.87 66.61 70.86 45.93 68.27 32.94 24.0 51.93 w. OWL 53.07 50.12 46.33 28.29 32.58 19.28 13.2 34.70 AlphaPruning 53.79 51.93 43.21 27.78 35.52 19.37 13.2 34.97 w. NeuronAl 52.35 51.93 43.33 28.08 32.28 18.52 12.0 34.07 Wanda 54.87 66.06 71.13 44.48 67.05 32.00 21.60 51.03 w. DSnoT 54.15 65.59 70.43 44.5 66.88 31.40 20.60 50.51 w. OWL 57.04 67.17 73.85 45.66 67.89 32.59 23.20 52.49 AlphaPruning 57.4 66.54 64.98 44.02 65.32 30.03 20.8 49.87 w. NeuronAl 557.04 66.06 70.34 46.37 68.6 33.02 23.2 52.09 OPT-6.7B multiflow 52.71 58.25 62.69 42.24 56.52 25.68 22.20 45.76 w. DSnoT 53.07 58.48 62.57 42.13 57.79 23.98 22.20 45.75 w. OWL 53.07 57.46 63.21 42.98 56.44 25.68 24.20 46.15 AlphaPruning 58.12 58.64 62.29 40.83 51.68 24.49 19.0 45.01 w. NeuronAl 52.71 58.33 62.32 42.19 56.4 25.0 21.6 45.51 Wanda 52.71 59.67 62.29 42.80 58.00 25.60 22.6 46.24 w. DSnoT 52.71 58.17 62.11 41.99 57.41 25.43 22.4 45.75 w. OWL 52.71 58.72 62.69 42.14 58.33 25.17 22.6 46.05 AlphaPruning 55.23 58.09 62.17 39.71 51.85 24.83 19.6 44.50 w. NeuronAl 52.71 61.25 62.39 41.95 57.28 25.09 22.0 46.10
Table 12: Accuracy on the seven Zero-Shot Tasks, computed over five different LLMs for three different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on two pruning algorithms (multiflow and Wanda) at 70% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to the uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B multiflow 53.79 50.43 53.49 28.43 46.59 20.48 15.6 38.40 w. DSnoT 52.71 54.06 54.92 28.38 44.15 21.5 15.6 38.76 w. OWL 52.35 55.33 61.99 30.31 54.59 24.23 18.4 42.46 w. AlphaPruning 47.29 51.3 37.71 25.85 26.64 20.9 13.0 31.81 w. NeuronAl 53.07 58.64 62.17 33.45 52.19 25.34 17.4 43.18 Wanda 52.35 53.2 62.14 28.31 44.87 20.99 17.4 39.89 w. DSnoT 52.35 51.54 60.98 28.33 41.62 21.08 15.4 38.76 w. OWL 52.71 53.59 62.05 30.09 48.61 22.53 18.8 41.20 w. AlphaPruning 47.29 49.88 60.86 25.52 25.55 22.01 14.0 35.02 w. NeuronAl 52.71 56.35 62.17 32.47 49.62 22.95 16.8 41.87 Llama1 7B multiflow 55.96 52.57 61.96 29.77 34.64 19.45 15.2 38.51 w. DSnoT 54.15 50.43 59.33 29.33 36.45 19.28 13.6 37.51 w. OWL 52.35 58.64 62.63 36.74 47.43 26.62 18.2 43.23 w. AlphaPruning 55.6 63.54 64.4 36.37 43.22 26.79 18.6 44.07 w. NeuronAl 57.76 61.72 63.3 38.19 50.88 26.88 22.8 45.93 Wanda 55.23 52.8 57.46 28.84 32.2 18.0 13.8 36.9 w. DSnoT 54.15 51.22 54.56 28.97 33.08 18.26 13.6 36.26 w. OWL 58.48 58.56 62.60 34.74 47.35 24.06 17.4 43.31 w. AlphaPruning 56.68 63.14 63.85 35.78 44.82 26.71 17.6 44.08 w. NeuronAl 55.96 59.27 63.12 36.94 50.0 26.11 20.6 44.57 Llama2 7B multiflow 52.71 50.99 62.05 28.52 33.04 17.92 13.6 36.98 w. DSnoT 52.71 50.99 59.72 27.92 32.58 16.81 13.0 36.25 w. OWL 52.71 56.12 62.05 32.40 42.42 19.88 18.6 40.6 w. AlphaPruning 52.71 58.64 62.2 35.17 41.33 22.87 17.6 41.5 w. NeuronAl 53.43 58.09 62.35 35.26 48.32 22.44 20.2 42.87 Wanda 52.71 48.46 49.94 28.09 30.39 19.2 11.8 34.37 w. DSnoT 52.71 50.36 47.77 27.67 30.6 17.32 12.2 34.09 w. OWL 52.71 55.96 62.11 31.86 43.73 20.65 17.0 40.57 w. AlphaPruning 52.71 61.33 62.2 34.82 43.43 22.1 17.6 42.03 w. NeuronAl 53.07 57.85 63.27 35.42 49.62 22.44 20.0 43.1 Mistral-7B multiflow 49.82 50.75 41.19 26.45 26.64 21.84 12.6 32.76 w. DSnoT 52.71 52.57 62.42 29.51 36.66 18.94 12.0 37.83 w. OWL 53.79 49.17 38.90 26.77 27.78 19.20 12.8 32.63 w. AlphaPruning 52.35 48.07 37.95 27.01 28.28 18.17 11.8 31.95 w. NeuronAl 52.71 50.75 38.29 27.16 28.75 17.75 13.0 32.63 Wanda 52.71 51.62 59.79 28.86 34.18 18.17 12.6 36.85 w. DSnoT 52.71 50.28 58.62 28.51 33.54 18.86 13.2 36.53 w. OWL 52.71 53.91 62.20 30.95 39.39 18.60 13.6 38.77 w. AlphaPruning 52.71 56.27 62.2 31.47 38.55 18.52 13.6 39.05 w. NeuronAl 52.71 60.62 62.17 34.8 44.28 20.31 16.0 41.56 OPT-6.7B multiflow 53.79 49.72 43.0 26.48 30.51 20.05 13.8 33.91 w. DSnoT 53.79 49.01 61.1 27.01 32.87 18.34 12.2 36.33 w. OWL 48.74 48.62 61.56 27.18 35.69 16.47 11.6 35.69 AlphaPruning 46.93 51.78 62.17 31.48 36.49 22.18 13.8 37.83 w. NeuronAl 50.54 50.99 62.17 31.3 40.74 22.27 17.4 39.34 Wanda 52.71 49.72 60.03 26.91 35.86 17.75 11.2 36.31 w. DSnoT 52.71 49.57 60.61 26.91 35.06 17.58 12.0 36.35 w. OWL 53.79 51.22 61.87 29.53 42.3 18.09 14.6 38.77 AlphaPruning 51.62 51.7 62.17 33.13 40.19 22.7 15.2 39.53 w. NeuronAl 50.90 51.07 62.17 30.54 40.78 21.16 15.4 38.86
Table 13: Accuracy on the seven Zero-Shot Tasks, computed over five different LLMs for three different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on two pruning algorithms (multiflow and Wanda) at 80% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to the uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B multiflow 52.71 50.12 51.04 26.15 27.44 19.37 13.4 34.32 w. DSnoT 50.18 49.49 37.83 26.32 27.82 19.97 13.8 32.2 w. OWL 53.07 51.22 56.76 26.30 30.85 18.52 15.2 35.99 AlphaPruning 44.77 48.78 37.83 25.74 26.18 20.31 11.4 30.72 w. NeuronAl 52.71 50.59 43.0 26.66 33.88 21.16 16.0 34.86 Wanda 53.07 49.25 62.17 25.93 27.44 20.99 14.6 36.21 w. DSnoT 53.07 50.99 38.04 26.17 27.10 20.73 13.0 32.73 w. OWL 52.35 51.22 60.21 26.51 29.88 20.05 13.6 36.26 AlphaPruning 43.32 49.8 38.29 25.94 25.67 21.16 11.0 30.74 w. NeuronAl 52.71 50.2 62.17 26.96 32.32 20.56 16.6 37.36 Llama-7B multiflow 47.29 50.91 40.03 26.17 26.77 21.16 12.4 32.1 w. DSnoT 46.57 50.43 37.83 26.02 27.06 20.05 12.2 31.45 w. OWL 50.18 50.04 45.47 26.74 27.65 20.48 12.2 33.25 AlphaPruning 56.68 49.96 62.17 26.44 27.23 21.25 11.2 36.42 w. NeuronAl 54.51 51.14 62.2 28.82 32.74 19.45 14.4 37.61 Wanda 47.29 49.88 37.83 26.34 26.47 20.99 12.8 31.66 w. DSnoT 46.93 50.36 37.83 26.03 26.56 21.33 13.4 31.78 w. OWL 47.29 49.88 37.83 26.67 27.19 19.54 11.6 31.43 AlphaPruning 52.35 49.17 61.19 26.58 26.77 20.73 10.6 35.34 w. NeuronAl 52.71 49.57 60.03 28.21 30.85 19.62 13.2 36.31 Llama2 -7B multiflow 53.43 48.86 37.83 26.35 27.48 21.25 13.2 32.63 w. DSnoT 52.71 48.86 37.86 26.17 26.60 20.39 13.0 32.23 w. OWL 52.71 49.49 37.83 26.62 26.94 19.11 12.6 32.19 AlphaPruning 52.71 50.67 38.01 26.62 26.77 20.14 11.2 32.30 w. NeuronAl 52.71 49.17 38.17 28.1 29.46 18.34 15.0 32.99 Wanda 47.65 49.41 37.83 25.82 26.52 20.82 14.6 31.81 w. DSnoT 53.07 47.91 37.86 26.09 27.23 20.73 13.0 32.27 w. OWL 52.71 50.83 37.83 26.52 27.27 19.37 12.8 32.48 AlphaPruning 52.71 48.54 37.89 26.42 27.44 19.45 12.2 32.09 w. NeuronAl 52.71 50.04 37.77 27.3 28.45 18.94 14.0 32.74 Mistral-7B multiflow 50.18 48.15 37.80 25.67 26.18 22.70 13.20 31.98 w. DSnoT 52.71 47.36 37.83 26.58 28.03 18.94 13.8 32.18 w. OWL 48.38 49.09 38.44 25.88 25.59 23.29 13.6 32.04 AlphaPruning 51.62 50.67 37.92 26.35 26.35 21.67 14.8 32.77 w. NeuronAl 52.71 48.46 37.83 26.07 27.57 19.03 14.6 32.32 Wanda 53.79 48.78 37.83 26.52 27.82 19.8 12.8 32.48 w. DSnoT 52.35 48.30 37.83 26.55 27.44 19.54 13.0 32.14 w. OWL 52.71 47.43 37.83 26.68 27.78 18.52 13.2 32.02 AlphaPruning 52.71 47.99 37.83 26.73 28.41 19.54 12.8 32.29 w. NeuronAl 52.71 51.14 38.04 27.28 28.66 20.31 13.4 33.08 OPT-6.7B multiflow 52.71 50.91 37.80 25.87 27.40 19.28 12.40 32.34 w. DSnoT 52.71 50.83 57.31 26.00 25.00 20.22 11.80 34.84 w. OWL 52.71 51.07 37.83 25.74 25.29 20.05 16.00 32.67 AlphaPruning 52.71 49.01 37.83 26.14 27.31 20.73 14.8 32.65 w. NeuronAl 53.07 50.12 62.23 26.44 31.48 20.48 12.6 36.63 Wanda 54.15 52.09 41.53 26.47 28.45 18.86 11.80 33.34 w. DSnoT 52.71 51.38 55.32 26.17 27.06 19.37 13.80 35.12 w. OWL 52.71 49.33 37.83 25.84 25.67 20.31 13.00 32.10 AlphaPruning 52.71 48.86 37.98 26.09 26.73 20.14 12.6 32.16 w. NeuronAl 52.71 50.67 41.83 26.43 30.05 18.69 11.0 33.05
A.1.3 Zero-Shot at 60%, 70%, and 80% sparsity with Magnitude
Table 14: Accuracy on the seven Zero-Shot Tasks, computed over five different LLMs for three different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on Magnitude at 60% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to the uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B Magnitude 57.04 62.83 51.38 42.56 66.33 35.41 28.2 49.11 w. DSnoT 54.51 64.09 42.32 41.09 66.25 34.98 26.6 47.12 w. OWL 55.23 62.59 48.81 42.53 67.38 38.48 28.4 49.06 w. AlphaPruning 57.76 56.2 47.34 35.16 58.46 33.28 23.2 44.49 w. NeuronAl 54.15 65.67 47.43 42.38 65.95 36.77 26.8 48.45 Llama-7B Magnitude 51.62 52.64 45.05 39.23 51.05 26.88 20.4 40.98 w. DSnoT 52.35 52.80 46.88 38.3 50.59 26.37 20.6 41.13 w. OWL 52.35 58.41 51.8 42.02 56.31 29.78 23.8 44.92 w. AlphaPruning 53.79 57.14 56.36 40.83 56.82 32.25 24.4 45.94 w. NeuronAl 50.54 56.04 57.46 40.63 55.26 29.86 24.4 44.88 Llama2 -7B Magnitude 51.26 55.8 41.19 36.97 50.17 26.96 16.2 39.79 w. DSnoT 53.79 56.04 42.87 38.3 53.28 27.9 19.8 41.71 w. OWL 51.99 57.3 46.15 42.56 56.65 30.46 19.4 43.50 w. AlphaPruning 52.35 61.88 58.1 46.13 58.59 31.83 24.8 47.67 w. NeuronAl 55.23 59.59 60.43 46.15 58.96 32.85 27.6 48.69 Mistral-7B Magnitude 55.23 62.19 66.36 48.74 67.05 33.19 22.6 50.77 w. DSnoT 55.6 62.35 68.53 48.28 67.51 33.11 23.2 51.23 w. OWL 53.79 64.48 72.17 49.39 68.14 33.87 23.8 52.23 w. AlphaPruning 54.87 63.93 74.89 47.24 63.89 32.0 23.0 51.40 w. NeuronAl 54.15 64.96 71.62 49.82 65.45 35.92 24.6 52.36 OPT-6.7B Magnitude 53.43 50.59 37.86 26.38 26.6 21.42 13.2 32.78 w. DSnoT 52.71 49.25 37.86 26.14 27.27 21.5 13.2 32.56 w. OWL 52.71 50.51 37.83 26.77 30.3 18.52 14.8 33.06 w. AlphaPruning 52.71 51.22 37.83 26.54 29.84 19.8 13.4 33.05 w. NeuronAl 52.71 53.91 39.11 33.23 37.71 24.06 16.8 36.79
Table 15: Accuracy on the seven Zero-Shot Tasks, computed over five different LLMs for three different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on Magnitude at 70% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to the uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B Magnitude 46.93 53.59 47.22 30.45 47.85 24.57 19.2 38.54 w. DSnoT 46.57 50.91 39.6 30.12 45.54 24.06 16.8 36.23 w. OWL 45.13 52.88 49.2 32.26 51.64 27.56 21.4 40.01 w. AlphaPruning 47.65 52.41 38.01 26.06 25.25 22.53 13.2 32.16 w. NeuronAl 47.65 53.51 52.81 33.22 53.87 30.03 20.6 41.67 Llama-7B Magnitude 53.43 49.96 37.92 27.59 31.73 22.44 16.6 34.24 w. DSnoT 52.71 51.7 37.83 27.71 30.26 22.7 15.4 34.04 w. OWL 53.07 51.38 38.38 33.14 39.31 24.15 16.8 36.6 w. AlphaPruning 52.71 53.04 39.48 36.18 43.56 24.83 21.2 38.71 w. NeuronAl 52.71 54.62 52.6 39.81 46.13 26.79 23.0 42.24 Llama2 -7B Magnitude 51.26 49.96 37.86 25.9 28.45 23.12 13.4 32.85 w. DSnoT 53.79 49.88 37.86 25.42 28.83 20.56 16.6 33.28 w. OWL 53.07 50.28 37.89 26.38 30.77 22.7 15.0 33.73 w. AlphaPruning 52.71 50.83 43.91 35.01 41.67 25.34 19.8 38.47 w. NeuronAl 54.51 55.41 64.86 33.25 42.09 27.47 21.2 42.68 Mistral-7B Magnitude 51.99 50.83 41.13 32.16 42.72 19.54 16.6 36.42 w. DSnoT 53.07 51.62 39.54 31.66 42.51 20.05 16.6 36.44 w. OWL 57.76 56.59 49.17 36.48 45.75 22.01 18.8 40.94 w. AlphaPruning 53.07 58.96 57.71 34.47 42.3 22.44 16.0 40.71 w. NeuronAl 53.79 58.56 62.6 38.6 44.23 26.28 21.0 43.58 OPT-6.7B Magnitude 52.71 49.8 37.83 25.88 26.68 21.33 12.4 32.38 w. DSnoT 52.71 49.96 37.83 25.87 27.19 20.14 13.6 32.47 w. OWL 52.71 50.59 37.83 25.81 25.46 21.25 12.8 32.35 w. AlphaPruning 52.71 52.01 37.83 26.21 27.74 20.9 13.0 32.91 w. NeuronAl 52.71 50.43 37.83 26.25 26.89 20.39 13.0 32.50
Table 16: Accuracy on the seven Zero-Shot Tasks, computed over five different LLMs for three different top-up pruning algorithms (DSnoT, OWL, and NeuronAl) on Magnitude at 80% sparsity. “Average” indicates the mean accuracy across tasks. The rows corresponding to the pruning algorithms refer to the uniform distribution.
Model Algorithm RTE WinoGrande BoolQ HellaSwag ARC-e ARC-c OBQA Average Phi-2.7B Magnitude 45.13 50.36 41.19 25.83 29.08 20.9 13.6 32.30 w. DSnoT 46.93 52.33 39.63 25.9 28.32 21.25 13.4 32.54 w. OWL 49.46 50.91 42.35 26.71 35.27 21.67 13.4 34.25 w. AlphaPruning 49.82 49.25 47.28 25.85 25.38 21.84 16.6 33.72 w. NeuronAl 50.54 52.25 42.69 26.21 28.7 22.1 12.8 33.61 Llama-7B Magnitude 46.21 49.96 53.98 25.69 24.83 21.84 13.8 33.76 w. DSnoT 52.35 51.85 38.47 25.52 26.39 21.42 16.0 33.14 w. OWL 48.38 48.93 44.74 25.76 26.35 21.08 15.8 33.01 w. AlphaPruning 51.26 50.2 39.54 26.25 28.28 20.99 15.8 33.19 w. NeuronAl 47.29 48.78 50.31 25.8 26.09 21.25 14.6 33.45 Llama2 -7B Magnitude 52.35 49.57 46.18 25.94 26.14 23.12 16.0 34.19 w. DSnoT 52.71 51.54 37.89 25.46 27.10 22.44 15.4 33.22 w. OWL 53.07 48.70 42.02 25.72 26.60 21.42 14.4 33.13 w. AlphaPruning 55.23 49.33 40.37 25.84 26.56 22.01 16.2 33.65 w. NeuronAl 54.15 50.36 52.69 26.82 29.67 20.14 14.6 35.49 Mistral-7B Magnitude 51.26 50.99 41.16 25.93 27.48 21.84 14.6 33.32 w. DSnoT 52.35 49.72 38.07 26.26 26.43 21.25 14.0 32.58 w. OWL 52.35 50.20 41.04 26.55 27.78 19.97 13.8 33.10 w. AlphaPruning 53.07 47.67 37.86 26.4 28.45 19.54 13.4 32.34 w. NeuronAl 53.43 50.04 61.04 28.56 32.2 22.53 16.2 37.71 OPT-6.7B Magnitude 52.71 49.49 37.83 25.79 26.39 21.25 13.0 32.35 w. DSnoT 52.71 49.57 37.83 25.78 25.63 20.65 12.8 32.14 w. OWL 52.71 49.80 37.83 26.05 26.73 21.16 13.2 32.50 w. AlphaPruning 52.71 49.09 37.83 25.84 26.14 20.82 12.8 32.18 w. NeuronAl 52.71 50.91 37.83 25.78 26.73 20.9 12.4 32.47
In Tables 14-16, we report the results of NeuronAl over Zero-Shot tasks using Magnitude pruning.
The results provided by NeuronAl turn out to be the best ones in 10 out of 15 cases while being the second best in 3 cases. It is also worth noticing that the performance gap between the Magnitude pruning and score-based pruning algorithms (such as Wanda or multiflow) is generally quite high. Hence, NeuronAl can improve the performance of Magnitude (in the standard setting with uniform distribution) only to a certain degree, since at high sparsity ratios (as the ones we test) the performance of Magnitude has been shown to be poor (Jaiswal et al., 2024).
clearpage
A.1.4 NeuronAl on LLama-13B at 70% sparsity
In Table 17, we present the results on LLama-13B at 60% sparsity. The results are in line with the ones at 70% and 80% sparsity presented in the main text since NeuronAl can outperform the competitors in 9 out of 15 cases while being second best in 4 cases.
Table 17: Perplexity of LLama-13B on the 3 Language Modeling datasets at 60% sparsity.
Algorithm Top-Up Llama-1 13B LLama-2 13B WikiText2 C4 PTB WikiText2 C4 PTB Magnitude Uniform 34.9 49.1 1413.7 10.1 13.3 457.5 DSnoT 33.6 41.3 604.8 10.1 13.3 376.7 OWL 28.5 36.6 255.0 8.9 11.6 217.2 AlphaPruning 29.7 37.2 207.6 8.2 11.0 172.8 NeuronAl 24.9 35.2 187.1 9.5 12.3 151.1 multiflow Uniform 8.7 10.9 66.5 16.3 21.0 211.6 DSnoT 8.4 10.8 58.7 8.3 11.3 217.9 OWL 7.9 10.0 48.1 8.5 11.2 120.3 AlphaPruning 7.9 10.3 47.8 8.6 11.3 118.1 NeuronAl 7.9 10.2 44.6 8.4 10.9 110.7 Wanda Uniform 8.8 11.2 72.1 8.4 11.5 146.0 DSnoT 8.5 11.0 66.4 8.3 11.4 131.3 OWL 7.6 9.8 47.6 7.5 10.2 98.0 AlphaPruning 7.7 10.1 45.4 7.6 10.3 99.7 NeuronAl 7.6 9.9 46.8 7.6 10.3 90.4
A.1.5 NeuronAl on SparseGPT at 60%, 70%, and 80% sparsity
In Tables 18-20, we present the results of NeuronAl on SparseGPT on the WikiText2, C4, and PTB datasets, using the block-only setting. To note: Since SparseGPT prunes and updates the weights from columns to rows, the row-wise step of NeuronAl cannot be included: indeed, it would force each row to have a different sparsity ratio, which is in contrast with the nature of SparseGPT.
Table 18: Perplexity on WikiText2 using SparseGPT.
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Uniform 15.8 10.4 10.2 9.4 13.4 OWL 15.8 9.4 9.2 9.1 14.2 NeuronAl 15.7 9.9 9.3 9.1 13.7 70% Uniform 28.9 27.3 27.3 22.0 20.5 OWL 27.7 19.2 20.5 18.6 21.6 NeuronAl 27.3 22.6 20.9 17.8 21.8 80% Uniform 131.0 207.0 122.1 98.4 95.7 OWL 107.5 93.8 84.3 77.2 80.8 NeuronAl 113.5 144.7 88.7 70.8 84.0
Table 19: Perplexity on C4 using SparseGPT.
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Uniform 19.0 12.8 12.9 13.0 15.3 OWL 19.2 11.7 11.6 12.4 15.8 NeuronAl 19.1 12.4 11.7 12.3 15.5 70% Uniform 28.6 28.3 31.5 27.8 22.4 OWL 28.2 21.1 22.8 23.7 22.4 NeuronAl 27.8 23.8 22.5 21.9 22.2 80% Uniform 98.7 136.2 104.8 86.5 72.5 OWL 79.7 68.3 73.4 66.2 65.4 NeuronAl 86.4 104.2 72.4 61.8 65.0
Table 20: Perplexity on PTB using SparseGPT.
Sparsity Top-Up Model Phi-2.7B LLama1 7B LLama2 7B Mistral-7B OPT-6.7B 60% Uniform 28.7 19.5 430.5 73.7 20.3 OWL 29.3 16.9 262.1 70.9 21.0 NeuronAl 28.2 18.2 249.2 67.2 20.6 70% Uniform 50.3 52.6 3780.0 153.2 32.0 OWL 51.0 37.0 1643.4 135.0 32.9 NeuronAl 47.3 40.5 861.6 123.4 32.8 80% Uniform 195.4 295.6 3201.7 316.2 102.3 OWL 141.4 162.3 5556.5 278.8 98.9 NeuronAl 156.7 260.2 3659.8 266.6 105.3
Using SparseGPT, the superiority of NeuronAl is less evident than with other pruning algorithms. Nevertheless, NeuronAl turns out to be the best-performing top-up algorithm in 5 out of 15, 8 out of 15, and 7 out of 15 cases, respectively for WikiText2, C4, and PTB. Interestingly, for lower sparsity, the gap between uniform and non-uniform distribution (both NeuronAl and OWL) is less remarkable than at higher sparsity. We explain these results with the inherent functioning of SparseGPT, which, differently from the other pruning algorithms, includes a weight reconstruction step. However, we can conclude that also in this case, our proposed approach turns out to be effective in many cases at increasing the task performance.